### v1

# Matching Hadronization and Perturbative Evolution: The Cluster Model in Light of Infrared Shower Cutoff Dependence

Abstract
In the context of Monte Carlo (MC) generators with parton showers that have next-to-leading-logarithmic (NLL) precision, the cutoff $Q_{0}$ terminating the shower evolution should be viewed as an infrared factorization scale so that parameters or non-perturbative effects of the MC generator may have a field theoretic interpretation with a controllable scheme dependence. This implies that the generator’s parton level should be carefully defined within QCD perturbation theory with subleading order precision. Furthermore, it entails that the shower cut $Q_{0}$ is not treated as one of the generator’s tuning parameters, but that the tuning can be carried out reliably for a range of $Q_{0}$ values and that the hadron level description is $Q_{0}$-invariant. This in turn imposes non-trival constraints on the behavior of the generator’s hadronization model, so that its parameters can adapt accordingly when the $Q_{0}$ value is changed. We investigate these features using the angular ordered parton shower and the cluster hadronization model implemented in the Herwig 7.2 MC generator focusing in particular on the $e^{+}e^{-}$ 2-jettiness distribution, where the shower is known to be NLL precise and where QCD factorization imposes stringent constraints on the hadronization corrections. We show that the Herwig default cluster hadronization model does not exhibit these features or consistency with QCD factorization with a satisfying precision. We design a modification of the cluster hadronization model, where some dynamical parton shower aspects are added that are missing in the default model. For this novel dynamical cluster hadronization model these features and consistency with QCD factorization are realized much more accurately.
preprint: UWThPh-2023-23 MCnet-24-05
1 Introduction
Multi-purpose Monte Carlo event generators (MCs) are indispensable tools to describe realistic, fully detailed hadronic final states for essentially all processes at collider experiments. Their underlying structure and components reflect the large hierarchies of energy scales involved in these processes. These scale hierarchies are also the basis of numerous factorization theorems in analytic QCD approaches, which are partly also reflected in the MC components. After the determination of hard scattering cross sections using matrix elements obtained from fixed-order perturbation theory, all-order leading contributions coming from the radiation of soft and collinear partons are resummed by parton shower algorithms. The latter evolve the incoming and outgoing energetic partons participating in the hard scattering to lower scales tied to an ordering variable. This evolution terminates when the ordering variable reaches a cut-off value, which we refer to as $Q_{0}$, and which defines a low scale kinematic restriction on the gluon radiation and gluon branching into quarks. The value of $Q_{0}$ is typically in the range between $0.5$ and $2$ GeV. Phenomenological models of hadronization then describe how hadrons are formed out of the partonic final states that have emerged from the parton shower. The parameters of these hadronization models are fixed through the tuning procedure which is based on a fit to a reference data set. In current state-of-the-art MCs it is common that the shower cutoff $Q_{0}$ is also treated as a hadronization parameter. This means that the value of $Q_{0}$ is fixed by demanding best agreement with the reference data. In collisions of extended objects, such as protons, also multi-parton interactions taking place in the initial stages of the collision are simulated. The modelling of multi-parton interactions is, however, not subject of the present paper.
One of the major recent advances in the development of MCs has been to include fixed-order (QCD) corrections for the description of the hard scattering process and the production of additional hard jets through NLO matching and multijet merging algorithms. These developments need to be complemented by more accurate parton shower algorithms to improve the description of the resummed higher order corrections arising from the soft and collinear dynamics. Such improvements for a large class of observables have only recently received more attention: While, since a long time, parton shower approaches such as coherent branching can be rigorously (i.e. analytically and numerically) proven to be accurate at the next-to-leading logarithmic (NLL) level for dijet-based $e^{+}e^{-}$ event-shape variables such as thrust or other event-shapes related to jet masses in the peak region Catani:1992ua ; Hoang:2018zrp ; Bewick:2019rbu ; Bewick:2021nhc , similar progress for dipole parton showers, which are more convenient for matching and merging, has only recently been achieved, see e.g. Dasgupta:2020fwr ; Forshaw:2020wrq ; Herren:2022jej for first steps in this direction. Going even beyond this level of accuracy for the parton shower evolution is also subject to an active area of research, but the implementation of these developments in the framework of multi-purpose MCs applicable for experimental analyses may likely still require significant time.
In the context of these advancements much less work has been invested in the developments of the MC hadronization models. However, the presence and impact of hadronization is an essential ingredient for a realistic description of infrared sensitive observables alongside with improved matching and merging scheme and NLL accurate parton shower algorithms. This is particularly important when MCs are employed in the context of the determination of scheme-dependent Lagrangian QCD parameters, where a systematic separation of perturbative and non-perturbative effects is crucial. This has for example been stressed since a long time for strong coupling determination analyses from $e^{+}e^{-}$ event-shapes, when non-perturbative hadronization corrections are determined from MC simulations Huston:2023ofk and the associated uncertainties are determined from simulations with different MC settings. The properties and quality of the MC hadronization models is also essential for the interpretation of the event generator’s top quark mass parameter $m_{t}^{\rm MC}$ determined in direct top mass measurements Hoang:2020iah ; CMS:2024irj , since they directly affect the top quark mass sensitive aspects of the observables used for the measurements Hoang:2018zrp .
Inspired by earlier studies of some of the authors in Ref. Proceedings:2018jsb , we argue that accurate and improved parton shower algorithms to be employed for MC predictions should not be considered independently of the MC hadronization models. In principle this is already obvious at the purely practical level in the context of the tuning procedure of an event generator to a set of reference data, since the tuning traditionally considers a combined fit of parton shower and hadronization model parameters. However, what we mean is, that eventually one should go beyond this practical level by demanding in addition that the combined matrix-element and parton shower partonic description, on one side, and the effects of the hadronization model on the other, by themselves have a well-defined field theoretic meaning in the context of QCD with a systematic and controllable scheme-dependence. It is obvious that this is particularly important when combining hadronization corrections extracted from MCs with high precision perturbative QCD predictions for the determination of hadron level eventshape distribution, or for the interpretation of $m_{t}^{\rm MC}$ obtained from the direct measurements. This implies that (i) the field theoretic meaning of the parton level description and the hadronization effects should not depend on any parameter subject to tuning hadronization effects and (ii) in particular that the parton shower infrared cutoff $Q_{0}$ should be considered as a factorization scale. This view has also been advocated in Ref. Platzer:2022jny , where the same reasoning has been applied for the construction of an algorithm for parton branching at the amplitude level Forshaw:2019ver based on an underlying factorization of jet cross sections. The way how the shower cutoff is implemented defines a particular scheme of this factorization, and the factorization property ensures that the hadron level description is scheme-invariant. This in particular implies that, at least within some limited $Q_{0}$ interval, the combined partonic and hadron-model prediction for infrared sensitive quantities should be independent of the $Q_{0}$ value. As it is probably too ambitious to demand exact $Q_{0}$-independence, in practice at least a systematic cancellation between the dominant linear $Q_{0}$-dependencies of the parton level predictions and of the associated hadronization corrections (at the precision level of the parton shower) should be realized. This $Q_{0}$-insensitivity with respect to linear $Q_{0}$ contributions is what we refer to as ’$Q_{0}$-independence’ in this article.
Within such a framework the analytic properties of a (NLO matched and) NLL accurate parton shower can be transferred in a fully controllable way to the hadron level description provided by the MC generator and new avenues to scrutinize the combined action of parton level and hadron level descriptions provided by the MCs are made possible. This level of control is also mandatory to study the impact of the shower cutoff $Q_{0}$ dependent top quark mass parameter $m_{Q}^{\rm CB}(Q_{0})$, which was recently proven to emerge from using the coherent branching algorithm for massive quark initiated $e^{+}e^{-}$ event-shape distributions Hoang:2018zrp , in the context of MC simulations for boosted top pair production.
For state-of-the-art MCs it is largely unknown to which extent their hadronization models satisfy the factorization criterium formulated above. It is obvious that the paradigm of parton-shower cutoff scale independence of the hadron level simulations imposes additional nontrivial constraints on the hadronization models. At least within some limited range for $Q_{0}$, they need the flexibility to match with the corresponding evolution of the parton shower description. Furthermore, the parton shower cutoff $Q_{0}$ should not be considered as a non-perturbative parameter, but always in a hierarchy $\Lambda_{\rm QCD}\ll Q_{0}\ll Q$, where $Q$ is the hard scale of the process of interest, and $\Lambda_{\rm QCD}$ is the intrinsic scale of QCD. Since typical parton shower cutoff values are around the charm mass scale, this interpretation is also practically feasible for current state-of-the-art MCs.
In this article we further promote the idea of the shower cutoff $Q_{0}$ being a factorization scale by providing an actual implementation of a hadronization model acting in this direction. This entails that the (linear) evolution in $Q_{0}$ of the parton level description and the corresponding hadronization effects individually follow the concrete predictions of QCD perturbation theory, yielding $Q_{0}$-independent hadron level descriptions in a controllable and systematic manner. To be definite, we carry out our considerations in the context of the angular ordered parton shower and cluster hadronization model as implemented in the Herwig 7 MC generator. Our new model will become available with an upcoming Herwig 7.4 release. We focus in particular on the $e^{+}e^{-}$ dijet event-shape distribution 2-jettiness. For the latter it is known since a long time ago Catani:1992ua ; Contopanagos:1996nh ; Almeida:2014uva that the coherent branching algorithm, which is the basis of the angular ordered Herwig parton shower, is NLL precise. For the 2-jettiness distribution the dominant linear shower cut $Q_{0}$-dependence of the parton shower description in the dijet limit can also be determined analytically at NLO (${\cal O}(\alpha_{s})$) and the Herwig 7.2 parton shower has been shown to be fully consistent with these analytic results also for quasi-collinear massive quarks in Ref. Hoang:2018zrp . Using analytic results for this linear $Q_{0}$-dependence and the association of the generator’s parton-to-hadron level migration matrix $T$ with the dijet shape function $S_{\rm had}$ appearing in QCD factorization and soft-collinear effective theory (SCET) for 2-jettiness in the dijet region Lee:2006nr , we can derive a QCD constraint concerning the migration matrix $T$, which can then be tested in detail. This constraint involves the linear $Q_{0}$-dependence of the first moment of the transfer matrix and demands in addition independence concerning the c.m. energy $Q$.
We find that the default cluster hadronization model of Herwig 7.2 does not satisfy this QCD constraint in a consistent manner.111 In this respect, Herwig 7.3, which has been released recently, behaves similar to Herwig 7.2. This motivates the construction of a novel and improved cluster hadronization model, which we call the “dynamical cluster model”. The dynamical cluster model satisfies the QCD constraints more accurately and leads to a significantly improved shower cut $Q_{0}$-independence of the hadron level description in the physically important shower cutoff interval $21~{}\mbox{GeV}<Q_{0}<$ 2 GeV. While even the behavior of the new dynamic hadronization model with respect to $Q_{0}$-variations is not yet perfect, the results of this article provide first important steps towards giving the hadronization effects provided by MC simulations a well-defined QCD meaning with a controllable scheme-dependence, and a stepping stone to achieve hadronization models fully consistent with QCD factorization, as outlined in Platzer:2022jny . Eventually variations of the shower cut may also be used as an instrument to quantity the theoretical uncertainty of event generator predictions in analogy to the canonical renormalization or factorization scale variations in analytic calculations. In this article we focus on hadronization in the presence of light quarks, which is a vital step towards understanding the interplay of the infrared cutoff, hadronization and the shower cutoff dependence of heavy quark mass parameters in an upcoming paper ourMCtopmasspaper .
The content of this article is a follows: In Sec. 2 we briefly review the basic conceptual components of the coherent branching algorithm implemented in the Herwig MC. As a novelty, we describe the access to the parton level involving quarks with current masses and massless gluons. This “true parton level” (which was already used in Ref. Hoang:2018zrp ) has not been directly accessible in earlier Herwig releases, but it is an essential prerequisite for the analyses in this work, and can also be compared to other approaches as those outlined in Reichelt:2021svh . In Sec. 3 we then recall the general aspects of the factorization between parton level and hadronization effect for MC simulations from Ref. Platzer:2022jny and the concrete resulting constraints on their (linear) shower cutoff $Q_{0}$-dependence for the 2-jettiness distribution in the dijet region obtained from the Herwig coherent branching implementation. These constaints, which are based on QCD factorization and Soft-Collinear Effective Theory (SCET), can be formulated in the form of RG evolution equations for distribution cumulants and moments and have been determined previously in Ref. Hoang:2018zrp . Here we also discuss the concrete relation of the non-perturbative shape function appearing in QCD factorization and the parton-to-hadron level migration matrix that can be extracted from the MC simulations which plays an essential role for our phenomenological analyses. The structure of the default cluster hadronization model available in the current Herwig release is explained in detail in Sec. 4 as a prequisite for the improved dynamical aspects of our novel cluster hadronization model introduced in Sec. 5. In this article, the consistency with respect to QCD factorization of the default and the novel dynamical hadronization models is tested through a number of analyses of the hadronization effects based on tunes for different fixed shower cutoff $Q_{0}$ values. These tuning analyses are described in detail in Sec. 6. Here we also address how to estimate uncertainties in the determination of the hadronization model tuning parameters, as this becomes relevant in the context of their cutoff scheme-dependence. Finally, in Sec. 7 we discuss the results of the $Q_{0}$-dependent tuning analyses from the phenomenological perspective, demonstrating a substantially better performance of the novel dynamical hadronization model with respect to the constraints imposed by QCD factorization. In Sec. 8 we conclude.
The reader not interested in the details concerning the Herwig hadronization models and our tuning analyses, may skip Secs. 4 and 5 and most of Sec. 6. However, we recommend to read Sec. 6.3 as it contains useful information for the understanding of the following phenomenological discussion.
2 Coherent Branching
2.1 General remarks
Coherent branching is a parton shower algorithm which is based on $1\to 2$ parton splitting processes incorporating QCD coherence through the ordering in an emission angular variable $\tilde{q}$, see e.g. the classic articles Catani:1990rr ; Catani:1992ua . In this work we use the implementation of coherent branching in the Herwig event generator, as previously described and analyzed in Gieseke:2003rz ; Hoang:2018zrp . For global observables in simple 2-jet processes such as event-shapes distributions in $e^{+}e^{-}\to$ hadrons, coherent branching is NLL precise. Here we do not discuss details of Herwig’s coherent branching algorithm [^2] We refer the reader to Ref. Hoang:2018zrp for a detailed discussion of features in Herwig’s coherent branching parton sbower relevant for the shower cutoff analyses in this article. but focus on its main features relevant for the kinematic properties of the emerging partonic final state and its characteristics relevant for the onset of the simulated hadronization dynamics taking place after the parton shower has terminated.
The first important aspect related to hadronization is that the coherent branching algorithm generates colour structures which are largely compatible with the space-time structure of hadronization in the sense that large-angle soft gluons will be emitted first in the parton shower evolution. Upon hadronization this effectively isolates the colors of the outgoing hard partons in the form of jets at the expense of forming a few, soft hadrons transverse to the hard jet momenta. The latter feature is also the basis of the analytic factorization methods for large-angle soft and collinear radiation dynamics such as in Soft-Collinear Effective Theory (SCET) Bauer:2002nz . The second important aspect is the preconfining property of the coherent QCD evolution. This property requires that, in the large-$N_{c}$ limit, color singlets (i.e. colour connected quark-antiquark pairs which form at the end of the shower through additional $g\to q\bar{q}$ branchings to be discussed in more detail in Sec. 4 and 5) acquire a universal invariant mass spectrum that is peaked at scales similar to the infrared cutoff $Q_{0}$ of the shower evolution and falls off exponentially for larger invariant masses. These ’clusters’, expressing the fact that color correlations are very local in phase space, can be interpreted as excited hadronic states and form the basic degree of freedom of the cluster hadronization model.
2.2 Kinematic Reconstruction, Reshuffling and True Parton Level
The coherent branching algorithm proceeds by generating, for each hard jet progenitor produced in the hard scattering or a decay process, a sequence of evolution variables: the angular scales $\tilde{q}$, momentum fractions $z$ and azimuthal orientations $\phi$ of the emissions. The values of these evolution variables are distributed according to the Sudakov densities and the splitting functions describing the individual branchings respecting the angular ordering restrictions already mentioned before Catani:1990rr . The full kinematics of the emerging partonic final state is, however, not yet determined during this process. Rather, it is inferred at the end of the shower evolution, when the sequence of these variables is terminated through an infrared resolution criterion. For the coherent branching algorithm this criterion is based on the transverse momentum of the branching (which is a function of $\tilde{q}$ and $z$) being larger than a cutoff value $Q_{0}$. Once the transverse momentum drops below $Q_{0}$, the parton shower evolution terminates. At this point a process of kinematic reconstruction takes place to determine a concrete physical final state with partons having four-momenta satisfying on-shell conditions and overall momentum conservation. This proceeds in two steps.
First, the momenta of the final state partons emerging from each progenitor are calculated from the sequence of evolution variables $\tilde{q},z,\phi$ according to the progenitor (forward) and partner (backward) direction associated to the evolving jet. At this point the four-momenta of all partons do not yet satisfy overall energy-momentum conservation since the total four-momenta of the partons emerging from each progenitor acquire invariant masses larger than the original progenitor masses. We call these invariant masses also progenitor virtualities below. Then, as the second step, a reshuffling algorithm is employed to balance the resulting jets against each other, maintaining overall energy-momentum conservation. In this algorithm the spatial momenta of all emitted final state partons in their common center-of-mass frame are rescaled by a global factor such that overall energy-momentum conservation is satisfied.
The (on-shell) masses of the final state partons emerging from the kinematic reconstruction procedure are in principle free parameters. In the context of processes involving only massless quarks and gluons, one would expect them to be zero. This is what we refer to as the true parton level, which can also be related to massless parton final states in standard computations in perturbative QCD. In general the true parton level is associated to the quarks having current quark masses $\hat{m}_{i}$. [^3] This true final state parton level has been employed and analyzed in detail already in our previous work Hoang:2018zrp . In practice, for a lght quarks a mass of $10$ keV is adopted. In that previous work we have also analyzed in depth how the parton shower algorithm affects the interpretation of the mass parameter for heavy quarks in relation to mass renormalization schemes. However, the initial steps to interface the parton level to the cluster hadronization model requires (much larger) constituent quark masses $m_{i}$ for all final state quarks. These constituent quark masses are parameters of the hadronization model and are constrained such that a cluster is kinematically allowed to decay at least into a pair of the lightest hadrons. Furthermore, the cluster hadronization model involves a branching of each gluon into $q\bar{q}$ pairs as the basis to form the initial clusters. The model thus assigns the gluons a mass $m_{g}$, such that this $g\to q\bar{q}$ splitting process is kinematically allowed at least for the lightest quarks with respect to their constituent masses. [^4] Below we also refer to the gluon mass $m_{g}$ as a constituent mass for simplicity.
For practical purposes, the previous default implementations of the parton shower in the Herwig event generator have only been providing a ’constituent’ parton level with the previously mentioned constituent quark and gluon masses. So the kinematic reconstruction and reshuffling procedures have been directly accounting for these masses, so that there has not been any direct access to the true parton level. In other words, the Herwig’s default parton level has already incorporated some aspects of hadronization. As long as the true parton level is of no relevance in the simulation, for example when the main focus is to describe experimental data, this is not an issue. However, for analyses (such as those carried out in this work) where the effects of hadronization need to be cleanly and accurately distinguished from the perturbative dynamics, it is. In particular, for events with large gluon multiplicities, the mass related kinematic effects can be quite significant given that tunes to data yield gluon mass values typically of the order of $1$ GeV.
In order to have access to the true parton level so that the effects of the hadronization model can be cleanly separated and quantified on an event-by-event basis, we have therefore extended Herwig’s functionality to perform the kinematic reconstruction and reshuffling (as described above) based on massless gluons and current quark masses $\hat{m}_{i}$. Subsequently, as the first part in the implementation of the hadronization model an additional reshuffling procedure is carried out which changes current quark masses $\hat{m}_{i}$ to constituent masses $m_{i}$ and provides the mass $m_{g}$ to the gluon. This reshuffling procedure is quite similar to the reshuffling carried out for the true parton level. First all final state partons are assigned their new mass ($m_{i}$ or $m_{g}$) and then again a global spatial momentum rescaling is carried out in the final state partons’ center-of-mass frame to ensure yet again overall energy-momentum conservation. The resulting partonic final state with quark constituent and gluon masses is not identical with, but very close to the constituent parton level of the default implementation. This procedure ensures that the process where the partons acquire quark constituent and gluon masses is separated completely from the shower evolution, so that it can be cleanly considered as a part of the hadronization model. This is vital to extract event-by-event migration matrices which describe how a parton level observable value correlates with a hadron level observable in a single event and enables us to calculate binned migration matrices for any observable by reading out two-dimensional histograms, see Sec. 6.3.
3 Expectations and Constraints on Hadronization
The default cluster hadronization model has predominantly been motivated by the preconfinement property of coherent QCD evolution, and was otherwise driven by minimal assumptions e.g. using only information on phase space and available quantum numbers as well as simple power laws for the dynamics within the model. [^5] A detailed description of the cluster model will follow in Sec. 4. The parton shower evolution provides kinematic and color connection information to the hadronization model that depends on the value of the scale $Q_{0}$, where the partonic evolution terminates. The value of the shower cut $Q_{0}$ has then typically been inferred through the tuning procedure, where also all parameters of the hadronization model (including also the quark constituent and gluon masses) are fixed in a fit to a set of reference data. This has assigned the shower cut effectively the role of an additional hadronization parameter even though its value affects the properties of the final states that emerge from the parton shower. Different hadronization models, while technically inter-operable among different types of parton showers and applicable for different values of the shower cut, have thus always shown an implicit dependence on $Q_{0}$ through the tuning procedure. However, this dependence has not been systemtically studied, neither to design and improve hadronization models, nor for systematic investigations of the uncertainty in the hadronization modeling or the size of the hadronization corrections. Nevertheless, it has been common practice to adopt hadronization effects extracted from MC simulations, through parton-to-hadron level migration matrices or by taking hadron-parton level ratios, as estimates for hadronization corrections for high-precision and potentially resummed perturbative QCD calculations, where for the latter the limit of a vanishing infrared regulator has been applied. A popular application of this kind is constituted by the previously mentioned strong coupling determinations from $e^{+}e^{-}$ event-shape distributions or jet rates ParticleDataGroup:2022pth .
In recent works Hoang:2018zrp ; Platzer:2022jny , however, we have been pointing out that the role of the parton shower cutoff $Q_{0}$ should be understood in the sense of an infrared factorization scale. This implies that the value of $Q_{0}$ and the way how the shower cut is implemented define a particular scheme how the partonic and the non-perturbative dynamics are separated and that the hadron level descriptions should exhibit some invariance under variations of $Q_{0}$. This factorization scale invariance would ensure that the hadronization model does not modify the infrared structure provided by perturbative QCD in an uncontrolled manner. In other words the shower evolution and the hadronization model should match at this scale, in the sense that the partonic final state provided by the parton shower at the scale $Q_{0}$ provides the starting point of the evolution of the hadronization model towards even lower scales where the non-perturbative dynamics at the scales of individual hadrons sets in. Thus at least within some limited range, the evolution of the hadronization model should be driven by the perturbative dynamics encoded also in the parton shower evolution. In Ref. Hoang:2018zrp we studied the parton shower evolution with $Q_{0}$ for Herwig’s coherent branching parton shower in detail for massless and massive quark $e^{+}e^{-}$ event shape distributions. We showed that the evolution is dominated by effects linear in $Q_{0}$ which can be quantified accurately through RG-evolution equations that can be calculated at NLO either from the coherent branching algorithm itself or from a common diagrammatic computation. This $R$-evolution equation is reviewed below in Sec. 3.2.
3.1 General Formulation
Schematically, our starting point can be summarized either as a factorisation at the level of the cross section or at the level of the colour density operator Platzer:2022jny , by studying a convolution of partonic and hadronic cross sections as
$\frac{{\rm d}\sigma_{H}}{{\rm d}w}=\sum_{n,m,c}\int\int{\rm d}\phi_{m}(p_{1},...,p_{m}|Q)\\ {\rm d}\sigma_{P,n,c}(q_{1},..,q_{n}|Q;Q_{0})S_{mn,c}(p_{1},...,p_{m}|q_{1},...,q_{n};Q_{0})\delta\left(w-W(p_{1},...,p_{m})\right)\,.$ (1)
Here ${\rm d}\phi_{m}(p_{1},...,p_{m}|Q)$ is the integration measure of the final-state hadron momenta and the term ${\rm d}\sigma_{P,n,c}(q_{1},..,q_{n}|Q;Q_{0})$ stands for the partonic cross section, including the partonic phase space integration of total momentum $Q$ and a certain colour-flow $c$. The term $S_{mn,c}(p_{1},...,p_{m}|q_{1},...,q_{n};Q_{0})$ represents the action of the hadronization model in converting $n$ partons into $m$ hadrons subject to a given model and momentum mapping inherent to it. $W(p_{1}...,p_{m})$ in turn is the observable’s definition at hadron level. In practice for an event generator we have $m>n$. We also stress the fact that such a probabilistic factorization is merely possible in presence of the large-$N$ limit and more generally would involve the presence of a colour flow in the parton level amplitude and its conjugate, see Platzer:2022jny for more details. In the above form it is clear that the demand of $Q_{0}$-independence implies sets of evolution equations for both the partonic and hadronic factor, which mix different partonic multiplicities, a generic feature which a hadronization model consistent with $Q_{0}$ independence must respect. To this end, we can use the evolution of the parton shower, which we can schematically write for a gluon exchange or emission with momentum $q_{n}$ in terms of a virtual $V_{c}(q_{1},..,q_{n})$ and real emission contributon $R_{c,c^{\prime}}(q_{1},..,q_{n})$ as
$\frac{\partial}{\partial Q_{0}}{\rm d}\sigma_{P,n,c}(q_{1},...,q_{n}|Q;Q_{0})=\int\left(V_{c}(q_{1},..,q_{n}){\rm d}\sigma_{P,n,c}(q_{1},...,q_{n}|Q;Q_{0})\right.\\ \left.-\sum_{c^{\prime}}R_{c,c^{\prime}}(q_{1},..,q_{n}){\rm d}\sigma_{P,n-1,c^{\prime}}(q_{1},...,q_{n-1}|Q;Q_{0})\right){\rm d}^{d}q_{n}\ ,$ (2)
where $V$ and $R$ include the definition of some momentum mapping, and can eventually be expressed in terms of splitting functions. Demanding that the hadronic cross section is independent of $Q_{0}$, one can then derive an evolution equation for $S_{mn}$ Platzer:2022jny . For the present purpose we note that Eq. (1) is an accurate analytic model of an event generator, and Eq. (2) one of parton shower evolution. We can also ask the question how such a factorization then gives rise to observable-specific hadronization corrections, which can be extracted in a MC sense by snapshots of events at the parton level and at the level of the final hadronic states. For $e^{+}e^{-}$ event-shapes this is discussed in the next subsection. Furthermore, we are naturally led to conclude that $Q_{0}$ should not be a tuning parameter and that we should investigate the $Q_{0}$ dependence of $S_{mn}$, for example, by studying the dependence of the tuned hadronization model parameters on variations of $Q_{0}$. As already mentioned in the introduction, the least we need to demand from an improved $S_{mn}$ is a smooth change across multiplicities in the sense that higher partonic multiplicities at lower $Q_{0}$ lead to comparable effect as smaller multiplicities at higher $Q_{0}$ do: the dynamics governing $S_{mn}$ near the infrared cutoff need to be the same as that of the shower close to the infrared cutoff. This puts severe constraints on the onset of the hadronization model, to be discussed in Sec. 4 and Sec. 5. Using the evolution equations implied for $S_{mn}$ for the construction of an evolving model is subject of additional ongoing work.
3.2 Thrust Distribution, Factorization and $R$-Evolution
The concrete observable we consider in most detail in this work is the 2-jettiness $\tau$ event-shape variable in $e^{+}e^{-}$ collisions, defined by
$\tau=\frac{1}{Q}\min_{\vec{n_{t}}}\sum_{i}(E_{i}-|\vec{n_{t}}\cdot\vec{p_{i}}|)\,,$ (3)
where $Q$ is the $e^{+}e^{-}$ c.m. energy, and the sum runs over all final state particles with momenta $\vec{p_{i}}$. The maximum defines the thrust axis $\vec{n_{t}}$. In the limit of vanishing hadron masses $\tau$ is identical to thrust. The difference is very small and does not play any role in the context of our studies. We therefore call $\tau$ also thrust sometimes in the rest of this article. Thrust is an IR safe and global shape variable and in the limit of small $\tau$, referred to as the dijet region, the events are characterized by two energetic back-to-back jets along the thrust axis. In this section we render factorization aspects of the coherent branching shower cutoff $Q_{0}$ and the resulting RG equation explicit for the thrust distribution in the dijet region. This RG equation serves as the basis of the concrete numerical studies we carry out in Secs. 7 for the default and our novel dynamical hadronization model.
Using the notations of Ref. Abbate:2010xh the hadron level thrust distribution in the dijet region can be written in the factorized form
$\frac{\mathrm{d}\sigma}{\mathrm{d}\tau}(\tau,Q)\,=\,$ $\int\limits_{0}^{Q\tau}\!\mathrm{d}\ell\;\frac{\mathrm{d}\hat{\sigma}}{\mathrm{d}\tau}\Big{(}\tau-\frac{\ell}{Q},Q\Big{)}\,\,S_{\rm had}(\ell)$ (4)
[…]
where $\mathrm{d}\hat{\sigma}/\mathrm{d}\hat{\tau}(\hat{\tau},Q)$ is the parton level distribution containing the resummed partonic QCD corrections. In the standard analytic QCD approach these computations are carried out in the limit of a vanishing IR cutoff, so that these perturbative QCD corrections encode terms of the form $\alpha_{s}^{n}\delta(\hat{\tau})$ and plus-distributions of the form $\alpha_{s}^{n}[\ln^{k}(\hat{\tau})/\hat{\tau}]_{+}$) to all orders of perturbation theory and potentially additional fixed-order corrections to improve the descriptions when $\tau$ increases. In this context $\mathrm{d}\hat{\sigma}/\mathrm{d}\hat{\tau}(\hat{\tau},Q)$ has been determined up to N3LL$+{\cal O}(\alpha_{s}^{3})$ order Becher:2008cf ; Abbate:2010xh . The exact form of the parton level distribution, which is based on an additional perturbative factorization between large-angle soft and energetic collinear radiation, is not relevant for the studies in this article. The relevant aspect is that the factorized form of Eq. (4) applies for any scheme to regulate the IR momenta in the partonic distribution $\mathrm{d}\hat{\sigma}/\mathrm{d}\hat{\tau}(\hat{\tau},Q)$.
The function $S_{\rm had}(\ell)$ is the shape function that describes the leading hadronization effects. For thrust in the dijet region it arises from the non-perturbative dynamics of the large-angle soft radiation in the vicinity of the hemisphere plane perpendicular to the thrust axis. Hadronization effects also exist for the energetic collinear radiation, but these are strongly suppressed and negligible. [^6] While large-angle soft dynamics is linearly sensitive to non-perturbative scales, the collinear dynamics is only quadratically sensitive and furthermore is associated to momentum fluctuations at higher scales of order $Q\sqrt{\tau}$ compared to the soft scales that are of order $Q\tau$. The shape function satisfies the normalization condition
[…]
It has an unambiguous definition related to vacuum-to-hadrons matrix elements of soft gluon Wilson lines. The scale $\ell$ is the non-perturbative light-cone momentum associated to coherent large-angle soft radiation respecting the thrust hemisphere constraint related to the thrust axis Korchemsky:1998ev ; Bauer:2008dt . The shape function is known to be universal for many event-shape distributions associated to the thrust axis hemisphere definition Lee:2006fn . In the canonical approach for analytic (and numerical) perturbative QCD computations the partonic distribution $\mathrm{d}\hat{\sigma}/\mathrm{d}\hat{\tau}(\hat{\tau},Q)$ and thus also the shape function are defined in the scheme of a vanishing IR cutoff. The shape function $S_{\rm had}(\ell)$ exhibits a peaked behavior for $\ell$ values around $1$ GeV and is strongly falling to zero for larger $\ell$. Due to the smearing effects induced by convolution over the shape function in Eq. (4) the peak of the hadron level $\tau$ distribution is shifted from the parton level threshold at $\hat{\tau}=0$ to positive values by an amount of order $\Lambda/Q$, where $\Lambda$ is in the range of $1$ to $2$ GeV. It is this peak region, which we focus on in our studies.
The point essential for our study is that the form of Eq. (4), which factorizes parton level and hadronization effects, is an umambiguous property of QCD independent of the scheme that is adopted to regulate IR momenta in the partonic distribution. In Ref. Hoang:2018zrp the NLL partonic thrust distribution was analyzed from the perspective of using the transverse momentum cut $Q_{0}$ that is employed in Herwig’s angular ordered parton shower to regulate IR momenta and keeping track of the dominant linear dependence on $Q_{0}$. [^7] In Ref. Hoang:2018zrp the analysis was carried out for the coherent branching algorithm as well in soft-collinear effective theory (SCET). The exact relation of NLL precision for the angular ordered parton shower and terms in the NLL$+{\cal O}(\alpha_{s})$ order counting in SCET was specified as well. It was found that only the large-angle soft radiation can cause a linear $Q_{0}$ sensitivity and that the partonic distribution $\mathrm{d}\hat{\sigma}/\mathrm{d}\hat{\tau}(\hat{\tau},Q,Q_{0})$ in the presence of a finite value of $Q_{0}$ can be written as
$\frac{\mathrm{d}\hat{\sigma}}{\mathrm{d}\hat{\tau}}(\hat{\tau},Q,Q_{0})\,=\,$ $\frac{\mathrm{d}\hat{\sigma}}{\mathrm{d}\hat{\tau}}\left(\hat{\tau},Q\right)+\frac{1}{Q}\Delta_{\rm soft}(Q_{0})\,\frac{\mathrm{d}^{2}\hat{\sigma}}{\mathrm{d}\hat{\tau}^{2}}\left(\hat{\tau},Q\right)$ (6)
[…]
with the gap function
$\Delta_{\rm soft}(Q_{0})\,=$  $\,16\,Q_{0}\,\frac{\alpha_{s}(Q_{0})C_{F}}{4\pi}\,+\,{\cal O}(\alpha_{s}^{2}(Q_{0}))$ (7)
and where $\mathrm{d}\hat{\sigma}/\mathrm{d}\hat{\tau}(\hat{\tau},Q)$ is the partonic thrust distribution for a vanishing IR cutoff shown in Eq. (4). Note that the simple form of Eq. (7) arises from a multipole expansion keeping the dominant linear dependence on $Q_{0}$ of the full result (which is also given in Ref. Hoang:2018zrp ). Upon convolution with the shape function this multipole expansion provides an excellent approximation to the full result. The strong coupling in the gap function $\Delta_{\rm soft}(Q_{0})$ is evaluated at the renormalization scale $\mu=Q_{0}$ since the gap function quantifies the effects of the unresolved (large-angle) soft radiation below the scale $Q_{0}$, which can only depend on the scale $Q_{0}$ in perturbation theory. The NLL precision of the coherent branching algorithm ensures that this $Q_{0}$ dependence is also realized by the Herwig simulation parton level results. The ${\cal O}(\alpha_{s}^{2})$ term indicated in Eq. (7) can only be specified once the cutoff prescription in the context of a more precise NNLL order shower evolution has been defined. Such a prescription is currently unknown and we therefore drop these higher order contributions from now on. The result for the gap function $\Delta_{\rm soft}(Q_{0})$ implies it satisfies the renormalization group equation
[…]
This evolution equation describes a linear scale dependence and has been called $R$-evolution in Refs. Hoang:2008fs ; Hoang:2008yj ; Hoang:2009yr .888 The point of the $R$-evolution is that the perturbation series for $\Delta_{\rm soft}(R)$ contains an ${\cal O}(\Lambda_{\rm QCD})$ IR renormalon, but that Eqs. (8) and (9) are renormalon-free. The NLL partonic thrust distribution at two different shower cutoff values $Q_{0}$ and $Q_{0}^{\prime}$ are therefore related by the equality
$\frac{\mathrm{d}\hat{\sigma}}{\mathrm{d}\hat{\tau}}(\hat{\tau},Q,Q_{0})\,=\,\frac{\mathrm{d}\hat{\sigma}}{\mathrm{d}\hat{\tau}}\left(\hat{\tau}+\frac{1}{Q}\Delta_{\rm soft}(Q_{0},Q_{0}^{\prime}),Q,Q_{0}^{\prime}\right)\,,$ (9)
where
$\Delta_{\rm soft}(Q_{0},Q_{0}^{\prime})\,=\,16\,\int\limits_{Q_{0}^{\prime}}^{Q_{0}}\mathrm{d}R\,\,\Bigl{[}\,\frac{\alpha_{s}(R)C_{F}}{4\pi}\,\Bigr{]}\,.$ (10)
We emphasize again that the NLL precision of the parton shower (for the thrust distribution) is essential, since otherwise Eqs. (6) and (9) and the evolution equation in Eq. (10), which can be computed in a straightforward way in analytic QCD computations, may not be realized by the parton level MC simulations.
Since factorization of the partonic thrust distribution and the non-perturbative shape function also applies in the context of a finite transverse momentum cutoff $Q_{0}$,
$\frac{\mathrm{d}\sigma}{\mathrm{d}\tau}(\tau,Q)\,=\,$ $\int\limits_{0}^{\tau}\!\mathrm{d}\hat{\tau}\;\frac{\mathrm{d}\hat{\sigma}}{\mathrm{d}\hat{\tau}}(\hat{\tau},Q,Q_{0})\,\,Q\,S_{\rm had}(Q(\tau-\hat{\tau}),Q_{0})\,,$ (11)
we can also derive the relation of the shape functions for two different cutoff values $Q_{0}$ and $Q_{0}^{\prime}$:
$S_{\rm had}(\ell,Q_{0}^{\prime})\,=\,S_{\rm had}(\ell-\Delta_{\rm soft}(Q_{0}^{\prime},Q_{0}),Q_{0})\,.$ (12)
Here we remind the reader that this relation is exact at the level of terms linear in the cutoff $Q_{0}$. Equalities (9) and (12) are concrete realizations of the more general and generic relations quoted in Sec. 3.1.
The quantity that conveniently quantifies the shape function’s linear dependence on $Q_{0}$ is the shape functions first moment [^9] The weight factor $1/2$ in the definition of the first moment is motivated by the fact that 2-jettiness is related to the sum of the two squared hemisphere (jet) masses in the small $\tau$ dijet region. The factor normalizes back to quantify the hadronization effects of a single jet.
$\Omega_{1}(Q_{0})\,\equiv\,\frac{1}{2}\int\!\mathrm{d}\ell\,\ell\,S_{\rm had}(\ell,Q_{0})\,,$ (13)
which leads to the relation
$\Omega_{1}(Q_{0}\prime)\,=\,\frac{1}{2}\,\Delta_{\rm soft}(Q_{0}^{\prime},Q_{0})+\Omega_{1}(Q_{0})\,.$ (14)
All other higher order cumulant moments do not have any linear cutoff dependence. The relation of the shape function’s first moment for different parton shower cutoff values shown in Eq. (14) is an essential prediction of QCD and plays an important role in our subsequent analysis of Herwig’s hadronization models.
3.3 Shower Cutoff Dependence for Thrust and Migration Matrix
At this point it is highly instructive to cross check the level of validity of relation (9) for the implementation of the angular-ordered parton shower in Herwig. As we already mentioned, Herwig’s angular-ordered parton shower is NLL precise for $e^{+}e^{-}$ event-shapes in the dijet limit. Even though the NLL precision is guaranteed conceptually, it is certainly useful to examine it for the practical implementation of the Herwig event generator. To this end, we consider the shower-cutoff dependence of the 2-jettiness cumulant
$\hat{\Sigma}(\hat{\tau},Q,Q_{0})\equiv\int\limits_{0}^{\hat{\tau}}\mathrm{d}\bar{\tau}\,\,\frac{\mathrm{d}\hat{\sigma}}{\mathrm{d}\bar{\tau}}(\bar{\tau},Q,Q_{0})\,.$ (15)
Using Eq. (9) it is straightforward to derive the partonic cumulant difference relation
$Q\,\frac{\hat{\Sigma}(\hat{\tau},Q,Q_{0})-\hat{\Sigma}(\hat{\tau},Q,Q_{0}^{\prime})}{\frac{\mathrm{d}\hat{\sigma}}{\mathrm{d}\hat{\tau}}(\hat{\tau},Q,Q_{0}^{\prime})}=\Delta_{\rm soft}(Q_{0},Q_{0}^{\prime})\,,$ (16)
where the normalization condition $\int\mathrm{d}\ell S_{\rm had}(\ell,Q_{0})=1$ applies. We make the important observation that the RHS of Eq. (16) is independent of $\hat{\tau}$ and the hard scale $Q$. As long as the soft-collinear approximations associated to the dijet region are valid, which are the basis of the factorization theorem in Eq. (4) and the shift relation in Eq. (6), this universality should hold to a good approximation. The range of thrust values where relation (16) is realized for Herwig’s angular ordered parton shower also indicates the expected range of validity of the factorization formula (11) and the relation of the shape function’s first moment for different cutoff values shown in Eq. (14) that tests consistency with QCD.
Figure 1: Binned cumulant differences of the parton level thrust distribution defined in Eq. (16) generated by the Herwig 7.2 angular ordered parton shower for $Q=45$ GeV (left panel), $91.2$ GeV (middle panel) and $200$ GeV (right panel) and shower cutoff values $Q_{0}=1$ GeV (blue) $1.5$ GeV (green) and $1.75$ GeV (red) with respect to the reference cutoff $Q_{0,{\rm ref}}=1.25$ GeV.
In Fig. 1 we have displayed the cumulant difference of Eq. (16) for $Q_{0}=1.0$ (blue), $1.25$ (orange), $1.5$ (green) and $1.75$ GeV (red) with respect to the reference result for $Q_{0}^{\prime}=1.25$ GeV for the hard scales $Q=45$, $91.2$ and $200$ GeV. We have employed bins of size $\Delta\hat{\tau}=0.02$, and each bin value is determined from the average of the cumulant at the respective upper and lower bin boundaries divided by the differential cross section of the bin. To visualize the impact of matching corrections which affect the 2-jettiness distribution outside the dijet region, four different Herwig matching settings have been used. The four different intensities for the same color correspond to (from darkest to the brightest color): (i) leading-order matrix element with matrix-element (ME) correction for hard QCD radiation (which is Herwig’s default); (ii) the same without ME correction; (iii) NLO matrix elements with multiplicative (POWHEG-type) matching and (iv) additive (MC@NLO-type) NLO matching, as available from the Matchbox framework Platzer:2011bc . The dotted horizontal lines correspond do $\Delta_{\rm soft}(Q_{0},Q_{0}^{\prime})$, where we use the strong coupling in the $\overline{\rm MS}$ scheme extracted from Herwig to solve the RG equation of Eq. (10). As expected, the impact of the matching and matrix element corrections increases for larger $\hat{\tau}$ values since the soft-collinear approximations in the coherent branching algorithms work more efficiently in the dijet region where $\hat{\tau}$ is small. We also see that the region in $\hat{\tau}$, where the matching corrections become sizeable, increases toward smaller $\hat{\tau}$ ranges for increasing c.m. energy $Q$. This indicates that the dijet 2-jettiness region around $\hat{\tau}=0$ decreases when the hard scattering scale gets larger and is consistent with the fact that the peak location and peak width scale with $\Lambda/Q$.
We see that, except for the very first $\hat{\tau}$ bin, where the no-radiation events and the implementation dependent details of the shower-cutoff dependence are still being resolved (and therefore not relevant for testing Eq. (16)), the partonic cumulant differences obtained from Herwig’s angular ordered shower are nicely compatible with the QCD value of $\Delta_{\rm soft}(Q_{0},Q_{0}^{\prime})$ for $\hat{\tau}$ up to about $0.2$. The visible discrepancies are related to quadratic and higher order effects in $Q_{0}$ which the linear approximation used in Ref. Hoang:2018zrp and expressed in Eq. (6) does not capture. Overall, we can conclude that the factorization formula (4) should be applicable for 2-jettiness values $\hat{\tau}$ up to around $0.2$ which well includes the peak region and a large fraction of 2-jettiness distribution tail.
Given that the angular ordered parton shower of Herwig exhibits the correct NLL shower cutoff $Q_{0}$ dependence, the hadronization model must have at an associated inverse $Q_{0}$ dependence so that the hadron level description is $Q_{0}$ independent and the shower cut $Q_{0}$ can be interpreted as an IR factorization scale. For the thrust distribution the effect of the hadronization in the MC generator appears in terms of a parton-to-hadron level migration matrix function $T$ connecting bins in the partonic thrust $\hat{\tau}$ to the hadron level thrust $\tau$. Writing the expression for binned distributions for simplicity in integral form, the generator’s hadron level thrust distribution can be written as
$\frac{\mathrm{d}\sigma}{\mathrm{d}\tau}(\tau,Q)\,=\,$ $\int\!\mathrm{d}\hat{\tau}\;\frac{\mathrm{d}\hat{\sigma}}{\mathrm{d}\hat{\tau}}(\hat{\tau},Q,Q_{0})\,\,T(\tau,\hat{\tau},\{Q,Q_{0}\}))\,.$ (17)
The transfer matrix $T$ is determined by reading out, for each event, the thrust values at the true parton level ($\hat{\tau}$) and at the hadron level ($\tau$) and by determining the resulting 2-dimensional histrogram. Unitarity ensures that $\int\mathrm{d}\tau\,T(\tau,\hat{\tau},\{Q,Q_{0}\})=1$ for each parton level $\hat{\tau}$. Comparing to the factorization analogue in Eq. (11) we see the close relation between the shape function $QS_{\rm had}(Q(\tau-\hat{\tau}),Q_{0})$, which satisfies the analogous normalization condition, and the migration matrix $T(\tau,\hat{\tau},\{Q,Q_{0}\})$. Changing variables to $k=Q\tau$ and $\hat{k}=Q\hat{\tau}$ we define the rescaled MC parton-to-hadron level migration matrix function
$\tilde{S}^{\rm MC}(k,\hat{k},\{Q,Q_{0}\})\equiv\frac{1}{Q}T\bigg{(}\frac{k}{Q},\frac{\hat{k}}{Q},\{Q,Q_{0}\}\bigg{)}\,.$ (18)
which is precisely the MC analogue to the shape function $S_{\rm had}(k-\hat{k},Q_{0})$.
To visualize the correspondence to the shape function $S_{\rm had}(\ell,Q_{0})$ more directly we can also consider the rescaled MC parton-to-hadron level migration function as a function of $\ell=k-\hat{k}$ for given values of $\hat{k}$, $Q$ and $Q_{0}$. We therefore define a shifted version of the rescaled migration matrix as a function of $\ell$,
$S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})\equiv\tilde{S}^{\rm MC}(\ell+\hat{k},\hat{k},\{Q,Q_{0}\})\,.$ (19)
Its first moment, defined in analogy to Eq. (13), should thus also satisfy the QCD constraint (14). Apart from the dependence on the shower cutoff scale $Q_{0}$, it is the additional potential dependence on the partonic momentum variable $\hat{k}$ and the hard scattering scale $Q$, which is not contained in the shape function $S_{\rm had}(\ell,Q_{0})$, that we will discuss in the phenomenological analyses in the later sections of this article. Details on how the hadronization model migration matrix function $\tilde{S}^{\rm MC}$ is extracted in Herwig are provided in Sec. 6.3. Examples for the migration matrix functions $\tilde{S}^{\rm MC}$ (left panels) and $S^{\rm MC}$ (right panels) are displayed in Fig. 10 for the default Herwig hadronization model and in Fig. 11 for the novel dynamical model.
The reader mostly interested in the phenomenology of the default and the novel dynamic hadronization models may now directly jump to Sec. 6.3.
4 The Default Cluster Hadronization Model
The cluster hadronization model is motivated by the preconfinement property of coherent QCD cascades. As we already briefly mentioned in Sec. 2, in the first step, gluons are split into quark-antiquark pairs such that (in the large-$N_{c}$ limit), colour-neutral $q\bar{q}$ systems, the ’clusters’, emerge. These clusters are interpreted as highly excited hadronic systems, and successively fission into lighter clusters, which, once below a certain threshold, decay into pairs of hadrons. Within our novel dynamical hadronization model the implementation of this final hadron decay process is the same as for the default model. So here we are mainly concerned with the gluon splitting, cluster formation and cluster fission processes, which we briefly review in this section focusing on the default Herwig implementation. These three processes are the important steps relevant for the matching to the $Q_{0}$-dependent infrared regime of the parton shower. For many of the details including a comprehensive description of the default implementation, we refer the reader to Ref. Bahr:2008pv .
4.1 Low-scale Gluon Splitting and Cluster Formation
After the parton shower has terminated, the final state consists of quarks and gluons. All gluons present thus need to undergo a branching into quark-antiquark pairs such that the colour neutral (mesonic quark-antiquark) clusters can be determined by the colour connections which the parton shower has produced. In the Herwig default hadronization model every gluon is assigned the same fixed constituent mass $m_{g}$, while for the novel dynamical model this gluon mass is generated dynamically, as explained in Sec. 5.1. As already mentioned in Sec. 2.2, in the existing default hadronization model implementation the gluon mass $m_{g}$ and the constituent quark masses $m_{i}$ have been implemented by a reshuffling procedure directly after the kinematic reconstruction, so that the true parton level (with current quarks and massless gluons) has not been accessible. For the hadronization model implementations used in this article we added access to the true parton level by first implementing current quarks and massless gluons (within the kinematic reconstruction and reshuffling at the end of the parton shower) and a separate subsequent reshuffling to constituent quarks and massive gluons. The latter ’constituent’ parton level constitutes the first step of the cluster hadronization model.
The gluon constituent mass $m_{g}$ of the default hadronization model is one of its parameters subject to the tuning procedure. In principle the same applies to the quark constituent masses $m_{i}$. However, for the cluster model they are highly constrained such that enough energy is available for a cluster to produce at least the lightest pairs of hadrons in its final decay. In practice the quark constituent quark masses are therefore fixed parameters of the default model, see Ref. Bahr:2008pv , as well as for our novel dynamical model.
After the reshuffling to the constituent parton level has been performed, each (now massive) gluon is forced to split into a light quark-antiquark pair. This decay is isotropic in the gluon rest frame, and the flavor of the emerging light quark-antiquark pairs are assigned randomly, see Sec. 7.1 in Ref. Bahr:2008pv for details. The associated probabilities are also tuning parameters, but they are not expected to carry any shower cutoff dependence and therefore fixed to the default in all our following analyses.
4.2 Cluster Fission
After the forced gluon splitting, the final state consists only of quarks and antiquarks having constituent quark masses $m_{i}$. The color connected quark-antiquark pairs are now combined into the clusters. For each cluster we have full information about its flavor content and the 4-momenta of its two constituents, which define the cluster’s mass $M$. The step that now follows in the hadronization is the cluster fission. Each cluster that fulfills the relation
${M}^{\rm{Cl}_{\rm{pow}}}\geq{\rm{Cl}_{\rm{max}}}^{\rm{Cl}_{\rm{pow}}}+(m_{1}+m_{2})^{\rm{Cl}_{\rm{pow}}}\,,$ (20)
where the $m_{i}$ are the masses of the clusters’s constituents, is considered “heavy” and will undergo fission. Otherwise it is called “light”. Cluster fission is a $1\to 2$ process, where one parent cluster is split into two daughter clusters. To do so a $q\bar{q}$ pair is popped from the vacuum, and together with the already two existing constituent quarks form two new color singlet clusters. If a daughter cluster is again “heavy” according to Eq. (20), it will itself undergo another fission, and so on, until all clusters are “light”. These final light clusters then decay into the pair of hadrons we already mentioned in Sec. 2.2. Depending on the type of cluster fission implementation (see below) and on the values of the parameters $\rm{Cl}_{\rm{max}}$ and $\rm{Cl}_{\rm{pow}}$, it can happen that also a cluster that is considered “heavy” according to Eq. (20) cannot undergo fission anymore because it is impossible to produce two physical daughter clusters which would be able to decay into hadrons individually. Also these clusters then decay directly into hadrons. The parameters $\rm{Cl}_{\rm{max}}$ and $\rm{Cl}_{\rm{pow}}$ are tuning parameters that govern how long the cluster fission proceeds and how heavy the clusters can be when they finally decay into hadrons. There are separate $\rm{Cl}_{\rm{max}}$ and $\rm{Cl}_{\rm{pow}}$ parameters for clusters containing charm and bottom quarks, but we have identified them in our analysis. The cluster fission condition (20) and the parameters $\rm{Cl}_{\rm{max}}$ and $\rm{Cl}_{\rm{pow}}$ are implemented for the default and our novel dynamic hadronization model. The difference is the dynamics of how the $q\bar{q}$ pair is produced from the vacuum.
In the default cluster fission the process is entirely one-dimensional in the sense that the momenta of the produced $q\bar{q}$ pair are directed along the axis defined by the cluster’s constituent 3-momenta in the cluster rest frame prior to the fission and that the direction of the 3-momenta of the original constituents remains unchanged. The only free parameters in this simplistic process are the masses of the two daughter clusters, $M_{1}$ and $M_{2}$. Only light flavored $q\bar{q}$ pairs can be popped from the vacuum and their flavor is picked randomly, with constant probabilities which are tuning parameters. The daughter cluster masses $M_{1}$ and $M_{2}$ are picked from a probability distribution that is a function of the parent’s mass $M$, the masses of the original constituents $m_{1}$ and $m_{2}$, the mass $m_{q}$ of the quarks popped from the vacuum and one or more tuning parameters $\vec{\lambda}$:
$\frac{\mathrm{d}^{2}P}{\mathrm{d}M_{1}\mathrm{d}M_{2}}=f\bigl{(}M,m_{1},m_{2},m_{q},\vec{\lambda}\bigr{)}\,.$ (21)
These general properties of the cluster fission are very similar for the default and the novel dynamical model. They differ in the way how the probability distribution in Eq. (21) is obtained. We note that these probability distributions are independent components and in principle not tied to how the forced gluon splitting discussed in Sec. 4.1 is handled. In order to better understand the novel aspects in the cluster fission of the dynamical model, we now describe briefly how the probability function is determined for the default cluster fission.
For the default fission the daughter cluster masses $M_{i}$ for each fission process are generated from the equation
$M_{i}=m_{i}+(M-m_{i}-m_{q})\times r_{i}^{1/\rm{PSplit}}\,\qquad i=1,2\,.$ (22)
Here, $r_{i}$ is a random variable drawn from a uniform distribution $\rm{unif}(0,1)$. Additionally, it is also required that the kinematic constraints
$M_{i}\geq m_{i}+m_{q}\,,\qquad M_{1}+M_{2}\leq M\,.$ (23)
are fulfilled. The parameter $\mathtt{PSplit}$ is the only tuning parameter for the cluster mass distribution of the default fission, i.e. $\vec{\lambda}=(\mathtt{PSplit})$. There are separate parameters for clusters containing charm and bottom quarks. The resulting double $M_{1}\leftrightarrow M_{2}$ symmetric differential mass distribution generated from Eqs. (22) and (23) reads
$\frac{\mathrm{d}^{2}P}{\mathrm{d}M_{1}\mathrm{d}M_{2}}\propto\Theta\bigl{(}M_{0}-M_{1}-M_{2}\bigr{)}\times g(M_{1})\times g(M_{2})\,,$ (24)
with
$g(M_{i})=\frac{\Theta\bigl{(}m_{i}+m_{q}<M_{i}<M_{0}-m_{q}\bigr{)}}{M_{0}-m_{1}-m_{q}}\biggl{[}\frac{M_{i}-m_{i}}{M_{0}-m_{q}-m_{i}}\biggr{]}^{-1+\rm{PSplit}}\,,$ (25)
and where we suppressed the normalization factor. Integrating let’s say over mass $M_{2}$ we can also write down the single differential mass distribution of the default cluster fission:
$\frac{\mathrm{d}P}{\mathrm{d}M_{1}}\propto$ $\frac{\Theta\bigl{(}m_{1}+m_{q}<M_{1}<M_{0}-m_{2}-m_{q}\bigr{)}}{M_{1}-m_{1}}\Biggl{[}\frac{M_{1}-m_{1}}{\Bigl{(}M_{0}-m_{1}-m_{q}\Bigr{)}\Bigl{(}M_{0}-m_{2}-mq\Bigr{)}}\Biggr{]}^{\rm{PSplit}}$
×$\qquad\times\biggl{[}\Bigl{(}M_{0}-M_{1}-m_{2}\Bigr{)}^{\rm{PSplit}}-m_{q}^{\rm{PSplit}}\biggr{]}\,,$ (26)
The smallest possible mass that a cluster generated in the fission can have is $m_{i}+m_{q}$. Therefore the lower bound on the mass of a cluster that can still undergo fission (i.e. being classified as “heavy” by Eq. (20)) is
$M_{\rm{min}}=m_{1}+m_{2}+2m_{q}\,.$ (27)
Neglecting for simplicity the quark constituent masses, the single differential cluster mass distribution in Eq. (4.2) reduces to the expression
$\frac{\mathrm{d}P}{\mathrm{d}M_{i}}\approx\frac{\Theta\bigl{(}M_{0}-M_{i}\bigr{)}}{M_{0}}\Bigl{(}\frac{M_{i}}{M_{0}}\Bigr{)}^{\rm{PSplit}-1}\Bigl{(}1-\frac{M_{i}}{M_{0}}\Bigr{)}^{\rm{PSplit}}\,.$ (28)
We remind the reader that the result in Eq. (28) provides the daughter cluster mass distribution for a single cluster fission process. The final resulting cluster mass distribution, at the point when the cluster fission processes of a single event have terminated, is more involved and also depends on the tuning parameters ${\rm{Cl}_{\rm{pow}}}$ and ${\rm{Cl}_{\rm{max}}}$ due to the heavy cluster condition in Eq. (20).
5 Dynamical Cluster Hadronization Model
Within Herwig’s cluster hadronization model implementation the two essential features that influence to which extend the hadronization model dynamics can properly match the infrared features of the parton shower are the gluon mass $m_{g}$, the kinematics of the subsequent forced gluon splitting in the first step of the hadronization process, and the dynamics behind the cluster fission. For a fixed gluon mass value the problematic aspect is that the gluon splitting process carried out by the parton shower entails in contrast a nontrivial distribution of invariant masses of the $g\to q\bar{q}$ process. This distribution depends on the value of shower cut $Q_{0}$. But it is also clear that a fixed gluon mass value can only mimic some averaged features of the parton shower splitting, and this prohibits an exact matching to the parton shower. For the cluster fission process, where an additional light quark-anti-quark pair is produced from the vacuum, the parton shower analogue is a radiated gluon which afterwords branches into the light quark-antiquark pair. For the default cluster fission process, the dynamical aspects encoded in these parton shower processes are missing. It can therefore not be expected that fixing the parameters $m_{g}$, ${\rm{Cl}}_{\rm{max}}$, ${\rm{Cl}}_{\rm{pow}}$ and $\rm{PSplit}$ from the tuning procedure should result in an exact matching to the parton shower.
The motivation for the construction of the dynamical cluster hadronization model is to add these parton shower aspects back to the model implementation such that the ultraviolet aspects of the hadronization modelling have an improved compatibility to the infrared behavior of the parton shower. The essential novel aspects are (i) a dynamically generated distribution for the gluon mass $m_{g}$ and parton-shower-like kinematics for the forced splitting in the initial stage of the hadronization and (ii) the implementation of a parton-shower-like dynamics behind the cluster fission process. These are explained in the following two subsections.
5.1 Dynamic Gluon Mass Distribution
The idea of the dynamic gluon mass distribution is to adopt essential features of the perturbative parton shower gluon splitting also for the non-perturbative gluon splitting in the hadronization model. In the parton shower, if a gluon branches into a quark-antiquark pair that will then be part of the final state (i.e. these quarks do not split any further and their momenta are set on-shell with current quark masses in the kinematic reconstruction process), the gluon adopts a finite virtuality associated to the quarks’ 4-momenta. The probability distribution of this gluon virtuality follows from the form of the splitting function and the implementation of the splitting algorithm.
For the construction a dynamical non-perturbative gluon splitting at scales lower than those probed by the parton shower, we implement important elements of this partonic branching by using the same splitting function with some modifications. The first obvious modification is that the splitting function implementation does not have the infrared $Q_{0}$-dependent cutoff of the parton shower and the second modification is to adopt a constituent mass $m_{q}$ for the produced quark pair (which automatically regulates the splitting function by the kinematic constraint $p_{g}^{2}>4m_{q}^{2}$). Furthermore, since we cannot use the perturbative QCD strong coupling $\alpha_{s}(\mu)$ for renormalization scales $\mu$ well below $1$ GeV, we adopt the frozen strong coupling value $\alpha_{s}(\mu_{0})$ at the scale $\mu_{0}=1$ GeV for scales $\mu<1$ GeV. As an additional feature we also account for the Sudakov form factor $\Delta(\tilde{Q}_{g}^{2},\tilde{q}^{2})$ to quantify a non-splitting probability between some scale $\tilde{Q}_{g}$ from which we start our “non-pert. shower” to the scale $\tilde{q}$ where the gluon splitting takes place. The scale $\tilde{Q}_{g}$ is a new tuning parameter of this gluon mass model (replacing the fixed gluon constituent mass $m_{g}$ appearing for the default version).
To keep things analytically trackable we approximate the Sudakov form factor with a Theta-function, because the scale hierarchies considered here are not large:
$\Delta(\tilde{Q}_{g}^{2},\tilde{q}^{2})\approx\Theta(\tilde{Q}_{g}^{2}-\tilde{q}^{2})\,.$ (29)
A similar approximation can also been used for the analytic Laplace space solution of the coherent branching algorithm in Refs. Catani:1992ua ; Hoang:2018zrp , see e.g. Eq. (4.11) in Ref. Hoang:2018zrp . Substituting $m_{g}^{2}=z(1-z)\tilde{q}^{2}$ for the gluon’s virtuality generated by the splitting, which we identify now with the gluon’s constituent mass $m_{g}$, the splitting probability for linear momentum fraction $z$ and gluon constituent mass $m_{g}$ for a single quark species is given by the expression
$\mathrm{d}P$ $\propto\Delta\Bigl{(}\tilde{Q}_{g}^{2},\frac{m_{g}^{2}}{z(1-z)}\Bigr{)}\,\mathrm{d}P_{g\to q\bar{q}}\Bigl{(}\frac{m_{g}^{2}}{z(1-z)},z,Q_{0}=0\Bigr{)}$
$=\frac{\mathrm{d}m_{g}^{2}}{m_{g}^{2}}\,\mathrm{d}z\,\frac{\alpha_{s}(m_{g}^{2})T_{F}}{2\pi}\Bigl{(}1-2z(1-z)+\frac{2m_{q}^{2}}{m_{g}^{2}}\Bigr{)}$
×$\hskip 28.45274pt\times\Theta(4m_{q}^{2}<m_{g}^{2}<z(1-z)\tilde{Q}_{g}^{2})\,\Theta\Bigl{(}z_{-}\bigl{(}\frac{m_{q}}{m_{g}}\bigr{)}<z<z_{+}\bigl{(}\frac{m_{q}}{m_{g}}\bigr{)}\Bigr{)}$ (30)
with
$z_{\pm}(x)\equiv\frac{1}{2}\Bigl{(}1\pm\sqrt{1-4x^{2}}\Bigr{)}\,.$ (31)
The restrictions on $z$ in the second line arise from the kinematics of the splitting process:
$0<p_{\perp}^{2}=z(1-z)m_{g}^{2}-m_{q}^{2}\quad\Rightarrow\quad 2m_{q}<m_{g}\,\,\mbox{and}\,\,z_{-}\bigl{(}\frac{m_{g}}{m_{q}}\bigr{)}<z<z_{+}\bigl{(}\frac{m_{g}}{m_{q}}\bigr{)}\,.$ (32)
Integrating over $z$ we can then obtain the probability distribution for the dynamic gluon mass,
$\frac{\mathrm{d}P}{\mathrm{d}m_{g}}\propto\frac{\alpha_{s}(m_{g}^{2})}{m_{g}}\,\biggl{[}$ $\Theta\Bigl{(}2m_{q}<m_{g}<\sqrt{m_{q}\tilde{Q}_{g}}\Bigr{)}\sqrt{1-\frac{4m_{q}^{2}}{m_{g}^{2}}}\Bigl{(}1+\frac{2m_{q}^{2}}{m_{g}^{2}}\Bigr{)}$
$+\Theta\Bigl{(}\sqrt{m_{q}\tilde{Q}_{g}}<m_{g}<\frac{\tilde{Q}_{g}}{2}\Bigr{)}\sqrt{1-\frac{4m_{g}^{2}}{\tilde{Q}_{g}^{2}}}\Bigl{(}1+\frac{3m_{q}^{2}}{m_{g}^{2}}-\frac{m_{g}^{2}}{\tilde{Q}_{g}^{2}}\Bigr{)}\biggr{]}\,.$ (33)
The resulting dynamic gluon mass distribution is shown as the red curve in Fig. 2 for the values $\tilde{Q}_{g}=6$ GeV (which is the typical value we obtain from the tuning analyses discussed later in Sec. 6) and the constituent quark mass $m_{q}=350$ MeV for the quark generated from the splitting. The gluon mass value for the standard tune of Herwig’s default forced gluon splitting, $m_{g,\rm{default}}=950$ MeV, is also indicated by the blue vertical line.
Figure 2: Dynamic gluon mass distribution for given $\tilde{Q}_{g}=6$ GeV and $m_{q}=350$ MeV (red curve) compared to the fixed default model gluon mass $m_{g,\rm{default}}=950$ MeV (blue vertical line).
For the implementation of the dynamic forced gluon splitting a value $m_{g}$ is drawn from the probability distribution in Eq. (5.1), that is then used for this gluon for the (second) reshuffling to constituent masses. After this reshuffling, each gluon with its dynamic gluon mass is now split into a light quark-antiquark pair in order to allow for clusters to be formed. While for the default implementation an isotropic quark-antiquark pair production process in the gluon rest frame has been used, for the novel dynamical model a process more closely resembling the parton shower dynamics is adopted. So the kinematics of the dynamical forced gluon splitting depends on the longitudinal momentum fraction $z$ and the azimuthal angle $\phi$ of the emission. While the latter azimuthal angle is uniformly distributed, the value $z$ is drawn from the distribution
$\frac{\mathrm{d}P}{\mathrm{d}z}$ $\propto\Bigl{(}1-2z(1-z)+\frac{2m_{q}^{2}}{m_{g}^{2}}\Bigr{)}\biggl{[}\Theta\bigl{(}2m_{q}<m_{g}<\sqrt{m_{q}\tilde{Q}_{g}}\bigr{)}\Theta\Bigl{(}z_{-}\bigl{(}\frac{m_{q}}{m_{g}}\bigr{)}<z<z_{+}\bigl{(}\frac{m_{q}}{m_{g}}\bigr{)}\Bigr{)}$
$\hskip 85.35826pt+\Theta\Bigl{(}\sqrt{m_{q}\tilde{Q}_{g}}<m_{g}<\frac{\tilde{Q}_{g}}{2}\Bigr{)}\Theta\Bigl{(}z_{-}\bigl{(}\frac{m_{g}}{\tilde{Q}_{g}}\bigr{)}<z<z_{+}\bigl{(}\frac{m_{g}}{\tilde{Q}_{g}}\bigr{)}\Bigr{)}\biggr{]}\,,$ (34)
which is just the $z$-dependent part of Eq. (5.1).
Using the standard decomposition into forward, backward and transverse momenta, the concrete expression for the quark momentum reads
$P_{q}^{\mu}=zP_{g}^{\mu}+\frac{m_{q}^{2}-(zP_{g}+q_{\perp})^{2}}{2z\,P_{g}\cdot\bar{n}}\,\bar{n}^{\mu}+q_{\perp}^{\mu}\,.$ (35)
The dependence on the azimuthal angle $\phi$ is contained in the transverse momentum component $q_{\perp}$ with respect to the axis $\bar{n}$, which has the concrete form
$q_{\perp}^{\mu}=\sqrt{-q_{\perp}^{2}}\Bigl{(}\cos(\phi)\,n_{\perp,1}^{\mu}+\sin(\phi)\,n_{\perp,2}^{\mu}\Bigr{)}\,,$ (36)
where
$P_{g}\cdot n_{\perp,i}=0\,,\quad\,\bar{n}\cdot n_{\perp,i}=0\,,\quad\,n_{\perp,1}\cdot n_{\perp,2}=0\,,\quad\,n_{\perp,i}^{2}=-1\,.$ (37)
Using in addition the relations
$P_{g}^{2}=m_{g}^{2}\,,\qquad-q_{\perp}^{2}=z(1-z)m_{g}^{2}-m_{q}^{2}\,,$ (38)
we finally arrive at
$P_{q}^{\mu}=zP_{g}^{\mu}+\frac{m_{g}^{2}(1-2z)-2P_{g}\cdot q_{\perp}}{2\,P_{g}\cdot\bar{n}}\,\bar{n}^{\mu}+q_{\perp}^{\mu}\,.$ (39)
For the antiquark momentum we have $P_{\bar{q}}^{\mu}=P_{g}^{\mu}-P_{q}^{\mu}$, which just corresponds to the replacements $z\to(1-z)$ and $q_{\perp}^{\mu}\to-q_{\perp}^{\mu}$.
At this point we still need to provide a concrete expression for the backwards light-like direction $\bar{n}$. In the parton shower the direction $\bar{n}$ is uniquely defined as the backwards direction to the momentum of the progenitor of the branching tree. However, for the forced gluon splitting this information is not available any more, and in principle there is no unique “correct” way assigning the direction of $\bar{n}$ so that we need to decide on a prescription. However, all choices should be close (or collinear) with the progenitor’s backward direction. To determine $\bar{n}$ we define a “new progenitor” momentum $P$, from which we obtain the light-like backwards direction as
$\bar{n}^{\mu}=\begin{pmatrix}1\\ \frac{-\vec{P}}{|\vec{P}|}\end{pmatrix}\,.$ (40)
To do this we first identify the momenta $P_{i}$ of the two (large-$N_{c}$) color connected final state partons of the splitted gluon, which either are a quark and antiquark, a quark and a gluon or an antiquark and a gluon. For our implementation we adopt one of these partons as the “new progenitor”, choosing the one whose direction of motion leads to a smaller transverse momentum of the gluon. With this choice for $\bar{n}$ together with the values for $m_{g}$, $z$ and $\phi$ we have now fully determined the quark momentum in Eq. (39).
Since the gluon splitting function is symmetric in $z$ around $z=1/2$, the quarks and antiquarks generated in the splitting have the same probability for both going in the direction of their respective color partner (with that they will then form a cluster) as for both going in the opposite direction of their color partner. There is in principle nothing wrong with that since this also happens in the parton shower. However, since one can argue that the non-perturbative splitting considered here is already the first step of the cluster formation, it is more “natural” that the quarks are predominantly emitted in the direction of their color partner. [^10] In fact, colour reconnection models Gieseke:2017clv ; Gieseke:2018gff would prefer to align the colour connections in such a way as to minimize the cluster masses. We therefore restrict the range of $z$ for the emitted quark to $z<1/2$ when the progenitor is the color connected quark and to $z>1/2$ when the progenitor is the antiquark.
So far we have discussed only one quark flavor of mass $m_{q}$ being produced in the splitting process. The gluon splitting, however, generates all three light flavors. In the default model implementation there have been fixed (tuned) probabilities $p_{i}$ for the three possible flavors $i=u,d,s$ to be chosen when the gluon is split. In the dynamic gluon splitting we go a different way. Writing the (not normalized) gluon mass distribution for light flavor $q$ in Eq. (5.1) as $\mathrm{d}P(m_{g},\tilde{Q}_{g},m_{q})$, the full gluon mass distribution $\mathrm{d}P(m_{g},\tilde{Q}_{g})$ is obtained from the sum over all light quark flavors
$\frac{\mathrm{d}P(m_{g},\tilde{Q}_{g})}{\mathrm{d}m_{g}}\propto\sum_{i=u,d,s}\frac{\mathrm{d}P(m_{g},\tilde{Q}_{g},m_{i})}{\mathrm{d}m_{g}}\,.$ (41)
We draw the gluon masses from this distribution and then do the reshuffling from the true to the constituent parton level. When coming do the point where we have to split the gluon, we have to decide for one light flavor. This is done randomly where the gluon mass dependent probability for flavor $i$ reads
$p_{i}=\frac{\mathrm{d}P(m_{g},\tilde{Q}_{g},m_{i})}{\mathrm{d}m_{g}}\times\biggl{(}\sum_{j=u,d,s}\frac{\mathrm{d}P(m_{g},\tilde{Q}_{g},m_{j})}{\mathrm{d}m_{g}}\biggr{)}^{-1}\,.$ (42)
This ensures that $\sum_{i=u,d,s}p_{i}=1$ and that $p_{i}=0$ if $2m_{i}>m_{g}$. We note that this implementation provides an exact treatment of the flavor-dependent gluon mass distribution $\mathrm{d}P(m_{g},\tilde{Q}_{g},m_{q})$ adapted to the basic setup of the cluster hadronization where the gluon splitting in the light quark-antiquark pairs takes place after the reshuffling to the constituent parton level.
5.2 Embedding Gluon Branching into Cluster Fission
Figure 3: Cluster formation and fission for the simple case of a quark-antiquark final state system produced by the hard scattering process. The upper path shows the process in the presence of a soft gluon that was just barely radiated perturbatively in the parton shower, which then splits into a quark-antiquark pair. The lower path shows the same system without any perturbative branching. After the formation of the first two primary clusters in the upper path, and the fission of the primary cluster into two secondary clusters in the lower path, the final state consists in both cases of two clusters. The further hadronization steps (more fission processes, cluster decay, hadron decays) are identical for the two paths.
The idea behind the novel dynamical cluster fission is, similar as for the dynamic gluon mass distribution, to implement a fission process that mimics important aspects of the parton shower dynamics. To illustrate our implementation let us have a look on the generic aspects of the process of cluster formation and cluster fission for the case of a simple quark-antiquark system, once in the presence of a (soft) gluon that was radiated in the parton shower and once without any perturbative radiation, see Fig. 3. If a gluon has been radiated by the parton shower (upper path in Fig. 3), in the first step of the hadronization model all particle momenta are reshuffled to their constituent masses and the gluon is forced to split into a quark-antiquark pair as we have discussed in Secs. 4.1 and 5.1. In the next step the color-connected quarks and antiquarks are combined into two clusters. At this point the final state consists of two clusters. If instead no gluon has been radiated in the parton shower (lower path in Fig. 3), we start the hadronization with only the quark-antiquark pair. In the first step of the hadronization, the two quarks’ momenta are reshuffled to their constituent masses, and subsequently combined into one single primary cluster. If this cluster is heavy, it undergoes cluster fission and splits into two lighter clusters. At this point the final state consists of two clusters as well.
Let us now assume that the perturbative radiation of the gluon in the shower in the upper path of Fig. 3 happens at a very low sale just slightly above the cutoff scale $Q_{0}$. So the upper path in Fig. 3 (perturbative gluon radiation, formation of two primary clusters) and the lower path in Fig. 3 (no perturbative radiation, with one primary cluster that undergoes cluster fission) are only separated by an very small shift of the cutoff $Q_{0}$ that either allows or vetoes the soft gluon radiation in the shower. In both cases one ends up with two separate clusters in the final state (that can then either decay directly to hadrons if they are light, or undergo further fission processes if they are heavy, and eventually decay into hadrons). We see that a smooth transition between the perturbative and the non-perturbative dynamics at the cutoff scale requires that the cluster fission in the lower path mimics the dynamics of the parton shower’s gluon radiation and gluon splitting. The idea behind the novel dynamical cluster fission model is therefore a generalization of the parton-shower-like gluon splitting dynamics we have adopted for the forced gluon splitting and dynamic gluon mass distribution described in Sec. 5.1 to a constituent quark $q\to qg$ (or antiquark $\bar{q}\to\bar{q}g$) branching as the basis of the cluster fission. In the following we will explain some technical details of the implementation.
The dynamical cluster fission starts from a primary cluster made of a color-connected quark-antiquark pair emerging from the forced gluon splitting of Sec. 5.1. We consider the cluster’s rest frame, where the quark and antiquark, which we call constituents, are back-to-back. It is now randomly chosen (with equal probability) from which of the two constituents the intermediate gluon is being radiated. Let us call the momentum of the constituent from which the gluon is radiated $P_{1}^{\mu}$, and the momentum of the other constituent $P_{2}^{\mu}$. Both momenta are on-shell with respect to their constituents’ masses, i.e. $P_{i}^{2}=m_{i}^{2}$. We now define the like-light backwards direction $\bar{n}^{\mu}$ to parametrize the branching process from the backward direction of radiating constituent, i.e.
$\bar{n}^{\mu}\equiv(1,-\vec{P}_{1}/|\vec{P}_{1}|)\,.$ (43)
The momenta $k^{\mu}$ of the (anti)quark and $g^{\mu}$ of the gluon that emerge from the branching are
$k^{\mu}$ $=zP_{1}^{\mu}+\frac{m_{1}^{2}+p_{\perp}^{2}-z^{2}m_{1}^{2}}{z(\bar{n}\cdot P_{1})}\times\frac{\bar{n}^{\mu}}{2}+q_{\perp}^{\mu}\,,$ (44)
$g^{\mu}$ $=(1-z)P_{1}^{\mu}+\frac{m_{g}^{2}+p_{\perp}^{2}-(1-z)^{2}m_{1}^{2}}{(1-z)(\bar{n}\cdot P_{1})}\times\frac{\bar{n}^{\mu}}{2}-q_{\perp}^{\mu}\,,$ (45)
where $p_{\perp}^{2}=-q_{\perp}^{2}$. The gluon mass $m_{g}$ is the virtuality the gluon acquires from its splitting into a $q^{\prime}\bar{q}^{\prime}$ pair following the algorithm already described in Sec. 5.1. The invariant mass the radiating constituent acquires due to this ’mini-shower’, i.e. the $q\to qg$ (or $\bar{q}\to\bar{q}g$) and the $g\to q^{\prime}\bar{q}^{\prime}$ branching, reads
$M_{1}^{2}=(k^{\mu}+g^{\mu})^{2}=\frac{m_{1}^{2}}{z}+\frac{m_{g}^{2}}{1-z}+\frac{p_{\perp}^{2}}{z(1-z)}\,.$ (46)
Expressing the perp momentum in terms of the evolution variable for the gluon emission branching
$p_{\perp}^{2}=(1-z)^{2}(z^{2}\tilde{q}^{2}-m_{1}^{2})\,,$ (47)
the invariant mass $M_{1}$ can also written as
$M_{1}^{2}=m_{1}^{2}+\frac{m_{g}^{2}}{1-z}+z(1-z)\tilde{q}^{2}\,.$ (48)
At this point we need to restore energy-momentum conservation and modify the 3-momentum of constituents $1$ and $2$. We thus solve the equation
$M_{\rm{cl}}=\sqrt{|\tilde{p}|^{2}+M_{1}^{2}}+\sqrt{|\tilde{p}|^{2}+m_{2}^{2}}\,,$ (49)
which yields
$\tilde{P}_{1}^{\mu}$ $=(\sqrt{|\tilde{p}|^{2}+M_{1}^{2}},\,|\tilde{p}|\times\vec{P}_{1}/|\vec{P}_{1}|)\,,$ (50)
$\tilde{P}_{2}^{\mu}$ $=(\sqrt{|\tilde{p}|^{2}+m_{2}^{2}},\,-|\tilde{p}|\times\vec{P}_{1}/|\vec{P}_{1}|)\,.$ (51)
with $|\tilde{p}|=\frac{1}{2M_{\rm{cl}}}\lambda^{1/2}(M_{\rm{cl}}^{2},M_{1}^{2},m_{2}^{2})$ for the new 3-momentum of the two constituents. The concrete expressions for the momenta of the quark (or antiquark) and the gluon emerging from the branching of constituent 1 read
$\tilde{k}^{\mu}$ $=z\tilde{P}_{1}^{\mu}+\frac{m_{1}^{2}+p_{\perp}^{2}-z^{2}M_{1}^{2}}{z(\bar{n}\cdot\tilde{P}_{1})}\times\frac{\bar{n}^{\mu}}{2}+q_{\perp}^{\mu}$
$=z\tilde{P}_{1}^{\mu}+\biggl{(}2(1-z)m_{1}^{2}-\frac{z}{1-z}m_{g}^{2}+(1-z)z(1-2z)\tilde{q}^{2}\biggr{)}\times\frac{\bar{n}^{\mu}}{2(\bar{n}\cdot\tilde{P}_{1})}+q_{\perp}^{\mu}\,,$ (52)
$\tilde{g}^{\mu}$ $=(1-z)\tilde{P}_{1}^{\mu}+\frac{m_{g}^{2}+p_{\perp}^{2}-(1-z)^{2}M_{1}^{2}}{(1-z)(\bar{n}\cdot\tilde{P}_{1})}\times\frac{\bar{n}^{\mu}}{2}-q_{\perp}^{\mu}$
$=(1-z)\tilde{P}_{1}^{\mu}-\biggl{(}2(1-z)m_{1}^{2}-\frac{z}{1-z}m_{g}^{2}+(1-z)z(1-2z)\tilde{q}^{2}\biggr{)}\times\frac{\bar{n}^{\mu}}{2(\bar{n}\cdot\tilde{P}_{1})}-q_{\perp}^{\mu}\,.$ (53)
The momenta $\tilde{k}^{\mu}$, $\tilde{g}^{\mu}$ and the momentum $\tilde{P}_{2}^{\mu}$ of constituent 2 are now parametrized in terms of the splitting variables $z$ and $\tilde{q}$ that are determined from the $P_{q\to qg}$ splitting function, and the gluon mass $m_{g}$ that we have to draw from the dynamic gluon mass distribution as described in Sec. 5.1. The splitting variables $\tilde{q}$ and $z$ are drawn from the probability distribution given by the $q\to qg$ splitting function (including a flat distribution of azimuthal angles)
$\mathrm{d}P_{q\to qg}\propto\frac{\mathrm{d}\tilde{q}^{2}}{\tilde{q}^{2}}\frac{\mathrm{d}z}{1-z}\alpha_{s}\Bigl{(}z^{2}(1-z)^{2}\tilde{q}^{2}\Bigr{)}\biggl{[}1+z^{2}-\frac{m_{1}^{2}}{z\tilde{q}^{2}}\biggr{]}\Theta(z^{2}\tilde{q}^{2}-m_{1}^{2})\,,$ (54)
where the quark mass $m_{1}$ should always be understood to be the constituent mass of the branching quark, times the Sudakov form factor, that we approximate again via a Theta-function,
$\Delta(\tilde{Q}_{q}^{2},\tilde{q}^{2})\approx\Theta(\tilde{Q}_{q}^{2}-\tilde{q}^{2})\,.$ (55)
The scale $\tilde{Q}_{q}$, which is the quark analogue of the scale $\tilde{Q}_{g}$ for the gluon splitting in Eq. (29), is one of the new tuning parameters of the dynamic model, and sets the starting scale of the shower evolution for the non-perturbative (anti)quark branching in the cluster fission process.
The gluon with momentum $\tilde{g}^{\mu}$ now splits into a $q^{\prime}\bar{q}^{\prime}$ pair following the description of the forced gluon splitting in Sec. 5.1. However, the backwards direction $\bar{n}^{\mu}$ is given in Eq. (43). Furthermore, the analogue of the scale $\tilde{Q}_{g}$ (shown in Eq. (29)) which here we denote by $\tilde{Q}_{g}^{(f)}$ is related to the splitting variables $z$ and $\tilde{q}$ by the angular ordering relation111111 We note that there is also the option to use $\tilde{Q}_{g}^{(f)}$ as an additional tuning parameter that sets the scale of the gluon splitting in the fission independent of the (anti)quark branching. We have, however, not used this option in our analyses.
$\tilde{Q}_{g}^{(f)}=(1-z)\tilde{q}\,.$ (56)
We emphasize that this means that the gluon splitting scale $\tilde{Q}_{g}$ in the forced gluon splitting and $\tilde{Q}_{g}^{(f)}$ in the cluster fission are unrelated, hence the superscript $(f)$ for ’fission’. Finally, the color-connected quarks and antiquarks (one being a constituent and the other being either $q^{\prime}$ or $\bar{q}^{\prime}$) are paired into the two new clusters.
Refer to captionRefer to caption
Figure 4: Distribution of the daughter cluster masses $M$ in the fission for a parent cluster with mass $M_{0}=91.2$ GeV (left panel) and $M_{0}=10$ GeV (right panel). The solid lines show the distribution for the dynamical cluster fission for $\tilde{Q}_{q}=5~{}\mbox{GeV}$ (red) and $10$ GeV (blue). The dashed lines show the default mass distribution for $\mathtt{PSplit}=0.5$ (red) and $1$ (blue)..
The resulting distribution for the daughter cluster masses $M$ are shown in Fig. 4 for the examples of a parent cluster mass $M_{0}=91.2$ GeV (left panel) and $M_{0}=10$ GeV (right panel). The solid lines show the novel dynamical mass distribution for $\tilde{Q}_{q}=5~{}\mbox{GeV}$ (red) and $10$ GeV (blue). The dashed colored lines show the mass distribution for the default cluster fission for $\mathtt{PSplit}=0.5$ (red) and $1$ (blue). The values for $\tilde{Q}$ and $\mathtt{PSplit}$ are within the typical ranges obtained in our tuning analyses. All constituent quark masses are set to $m_{q}=0.35\,\rm{GeV}$. We see that the novel dynamical cluster fission yields a substantially richer structure of the daughter cluster mass distribution than the default model, which is merely assuming a power law behavior supplemented by constraints from the phase space limits. In contrast, the dynamical model naturally implements phase space constraints through the kinematics of the fission process and therefore incorporates physical thresholds where both light, but also very asymmetric heavy/light cluster configurations are allowed. This can be seen from the peak structures exhibited by the solid lines. It is typical that two peaks appear. The lighter peak corresponds to the cluster that is formed from the constituent quark of the parent cluster from which the gluon was radiated and the color connected quark generated from the gluon splitting. Since the gluon and the quark pair emerging from its splitting are radiated predominantly into the direction of that constituent quark, this cluster is typically light. The heavier peak then corresponds to the cluster that contains the other constituent quark. Furthermore, in the dynamical model there is always a smooth and fast suppression in the limit of vanishing cluster masses which is more physical than the sharp cutoff resulting from the simply power law mass distribution of the default model.
6 Description of the Shower Cutoff Dependent Tuning Analyses
6.1 Tuning Procedure and Reference Tune
[…]
The tuning procedure is performed with the software library Apprentice Krishnamoorthy:2021nwv . For the determination of the reference tune we convert the experimental data bins, which are provided in form of Yoda files, to a reference data file by using the Apprentice Python script “app-datadirtojson”. The second input required by Apprentice are Herwig samples for the corresponding observable bins generated from the Rivet analysis code related to the experimental data evaluated at different values of the tuning parameters within the parameter ranges. The Apprentice script “app-build” determines a polynomial interpolation from these samples for both height $f_{\mathrm{MC,k}}(\{p_{i}\})$ and statistical error [^12] We have fixed a bug in Apprentice which led to wrong values in the error interpolation. $\Delta f_{\mathrm{MC,k}}(\{p_{i}\})$ of each bin $k$ separately. The tunes are then produced by using the Apprentice script “app-tune2” which performs a weighted $\chi^{2}$ minimization by evaluating these interpolations. The goodness of fit function that Apprentice minimizes is
$\mathrm{GOF}(\{p_{i}\})=\sum_{k}w_{k}^{2}\frac{(f_{\mathrm{MC,k}}(\{p_{i}\})-f_{\mathrm{ref,k}})^{2}}{(\Delta f_{\mathrm{MC,k}}(\{p_{i}\}))^{2}+(\Delta f_{\mathrm{ref,k}})^{2}},$ (57)
where $f_{\mathrm{ref,k}}$ and $\Delta f_{\mathrm{ref,k}}$ are the height and the corresponding statistical error of bin $k$ of the reference histogram, respectively. In our analysis the weights $w_{k}$ are chosen to be common for bins belonging to the same observable.
[…]
In order to have a means to cross check that stability of interpolation procedure we carry out two independent analyses for the “app-build” interpolations, one using cubic and one using quartic polynomial orders (which yields 120 and 330 polynomial coefficients, respectively, for seven parameters). We use the more precise quartic interpolation as our default interpolation and the less precise cubic one as a reference for the stability checks. For the cubic and quartic interpolations we generate $10^{5}$ and $10^{6}$ events per parameter space point, respectively, to obtain the binned distributions. These leads to statistical errors that are, for both interpolations, approximately of the same size as our estimate for the interpolation uncertainties which are described in more detail below. The latter can be kept small by providing a sufficiently large oversampling factor compared to the number of coefficients of the polynomial interpolation. The number of sampled parameter space points are (hadronization model: interpolation order, number): (dynamic: cubic, 493), (dynamic: quartic, 2257), (default: cubic, 484), (default: quartic, 5380). This corresponds to at least a four-fold oversampling.
The “app-tune2” minimization script is run with the following options: The minimization algorithm is set to the truncated Newton (TNC) algorithm, where the starting point for the minimization is obtained by taking the point with the minimum $\mathrm{GOF}$-function value out of 100 randomly sampled points. The full minimization is then repeated 10 times to reliably find the global minimum. After determining the reference tune, the “app-tune2” script also automatically saves prediction histograms for all observables in an output file called “predictions_tnc_100_10.yoda” [^13] We have fixed a bug in Apprentice which led to wrong values in this prediction file.. It obtains these predictions by evaluating the Apprentice interpolation at the parameter values given by the reference tune. We convert this Yoda file to a reference data file by using the “app-datadirtojson” script again. This standard reference file provides the reference data for the $Q_{0}$-dependent analysis.
6.2 Treatment of the Hadronization Model Parameters
Tuning Parameters
The default and the novel dynamical cluster hadronization models both feature a number of tuning parameters. In our tuning studies we do not consider all of them as floating parameters to be determined by the tuning fits. Rather we consider those as floating which are associated with the ’hard scale’ for the hadronization dynamics (i.e. they should show some correlation to the parton shower’s IR cutoff $Q_{0}$) and on which the event-shapes, jet rates and charged particle multiplicities that we account for in the reference data depend in a significant way. Furthermore, we also include parameter which are related to low-scale hadron-specific processes unrelated to that ’hard scale’ which, from the physical perspective, should be rather $Q_{0}$-independent. Overall, for the default and for the novel dynamical hadronization models each, 6 parameters are treated as floating parameters in the fits to the reference data. We also refer to them as $p_{i}$ ($i=1,\ldots 6$) below. In this section we discuss the concrete features of these parameters in more detail. All other parameters, which are common to both hadronization models, are set to their default values.
The only hadron-specific hadronization model parameters we consider as floating tuning parameters are $\mathtt{PwtSquark}$ and $\mathtt{PwtDIquark}$. They control for example the strangeness and Baryon production rates and thus directly impact the charged particle multiplicities. Since hadron production takes place after the cluster fission process, $\mathtt{PwtSquark}$ and $\mathtt{PwtDIquark}$ appear for the default as well as for the novel dynamical hadronization model. Among all the hadron-specific parameters of the hadronization model it is important to include at least these two as floating parameters in the tuning analyses to maintain a realistic correlation between various parameters that affect the charged particle multiplicities, but in the end also have an impact on other observables. We also note that several of the cluster fission parameters are flavour dependent, in order to allow them to account for effects of the heavy quark masses. Sensitivity to those can only be gained by explicitly considering flavour dependent observables, but otherwise their role is no different from their light-quark counterparts. Since we are not specifically addressing heavy quark fragmentation in our analysis, we choose to identify those heavy flavor specific parameters with their light quark counterparts for the studies in this article.
Apart from the two hadron-specific parameters $\mathtt{PwtSquark}$ and $\mathtt{PwtDIquark}$ just explained, there are two additional hadronization model parameters which we treat as floating tuning parameters and which are common to both hadronization models. These are the parameters $\mathtt{Cl}_{\mathtt{max}}$ and ${\mathtt{Cl}_{\mathtt{pow}}}$. They appear in the “heavy” cluster condition of Eq. (20) which determines whether the fission of a (heavy) cluster into two lighter clusters takes place or whether that cluster is already considered light and decays into hadrons, see Sec. 4.2. While $\mathtt{Cl}_{\mathtt{max}}$ (which has dimension of mass) sets the overall scale of the heavy cluster fission threshold, the dimension-less parameter ${\mathtt{Cl}_{\mathtt{pow}}}$ essentially quantifies a smearing of the threshold that also depends on the masses of the cluster’s constituents.
The two remaining hadronization parameters we consider in our $Q_{0}$-dependent tuning analyses differ for the two hadronization models. For the default hadronization model these are the gluon constituent mass $m_{g}$ and the dimension-less parameter $\mathtt{PSplit}$. The gluon constituent mass $m_{g}$ is important for the kinematics of the forced gluon splitting taking place at the initial states of the hadronization process, see Sec. 4.1. For the novel dynamical hadronization model the fixed gluon mass $m_{g}$ is replaced by a gluon mass distribution. The parameter $\mathtt{PSplit}$ governs the shape and steepness of the daughter cluster mass distribution according to Eq. (22) in the cluster fission algorithm described in Sec. 4.2. It has no counterpart in the novel dynamical model, where the daughter cluster mass distribution is generated from a splitting process. We refer to our discussion of Fig. 4. For the novel dynamical hadronization model the two remaining hadronization parameters are $\tilde{Q}_{g}$ and $\tilde{Q}_{q}$ both which have dimension of energy. The scale $\tilde{Q}_{g}$ is the ‘hard energy scale’ of the $g\to q\bar{q}$ branching process that is the basis of the forced gluon splitting in the dynamical model, see Sec. 5.1. The scale $\tilde{Q}_{q}$ is the ‘hard energy scale’ of the $q\to qg$ (or $\bar{q}\to\bar{q}g$) branching process that governs the cluster fission process in the dynamical model, see Sec. 5.2.
Since $\tilde{Q}_{g}$ and $\tilde{Q}_{q}$ represent the scales where the non-perturbative splitting and fission processes start we can expect some linear dependence of their fitted values on the value of the parton shower cutoff $Q_{0}$, if a proper matching of the dynamical hadronization to the parton shower is realized through the tuning to the reference data. Furthermore, in this case, the other four parameters of the dynamical hadronization model should be rather insensitive to the shower cutoff value since they govern dynamical aspects that are only of non-perturbative nature taking place at scales below the hard scales $\tilde{Q}_{g}$ and $\tilde{Q}_{q}$. On the other hand, for the default hadronization model, which is not designed to provide any systematic matching and where $Q_{0}$ essentially plays the role of just another hadronization parameter, such behavior cannot be naturally expected. To which extent these expectations are actually met by the outcome of our tuning analyses is subject to our phenomenological discussion in Sec. 7.2.
Error Estimate
The description up to this point is complete with regards to the determination of the central values $p_{\mathrm{i,cent}}$ ($i=1,\ldots 6$) of the hadronization tuning parameters. Interestingly, at this time there is no general canonical approach to estimate the uncertainties of MC hadronization model tuning parameters. However, in the context of having hadronization effects being defined in a particular scheme, it is also relevant to quantify an uncertainty on its parameters. In the following we explain the prescription we adopt for an uncertainty estimate of the tuning parameter we treat as floating in our tuning fits. As there is no canonical approach, our prescription is to some extent ad-hoc and only provides a first step towards are systematic treatment of the tuning parameter uncertainties. However, we believe that viewed together with the differences obtained from the cubic and quartic interpolations it provides a sufficiently fair treatment at this point.
There are two sources of uncertainties we consider. The first is the statistical uncertainty related to the number of events in our MC simulations. The second is an estimate for the “app-build” interpolation error. The statistical error is determined from the inverse of the Hessian matrix $H_{ij}\equiv\partial^{2}(\mathrm{GOF})/(\partial p_{i}\partial p_{j})$ of the fit at the best fit point multiplied with a heuristic rescaling factor associated to the effective degrees of freedom induced by the weights $w_{k}$:
$\Delta p_{i,\mathrm{stat}}(Q_{0})=\sqrt{2*(H^{-1})_{i,i}(Q_{0})*\left(\sum_{k}{w_{k}^{2}}\right)^{2}\left/\middle(\sum_{k}{w_{k}}\right)^{2}}\,.$ (58)
To estimate the interpolation uncertainty we analyze the difference for the tuning parameter obtained from the reference data (that is based on the Apprentice interpolation) versus data generated from a full Herwig simulation run, which we call exact Herwig data. Given the central values for the hadronization parameters ${p_{i,\mathrm{cent}}(Q_{0})}$ obtained at cutoff $Q_{0}$, we can carry out a second tuning fit for the cutoff $Q_{0}$ with the reference data being replaced by exact data obtained from a Herwig run using the tuning parameters $p_{i,\mathrm{cent}}(Q_{0})$. This yields a new set of best tuning parameter values which we call $p_{i,\mathrm{refH}[Q_{0}]}(Q_{0})$, where the subscript $\mathrm{refH}[Q_{0}]$ stands for the shower cutoff where the exact data is generated and the argument $(Q_{0})$ for the cutoff where the tuning fit is carried out. The difference between these two sets of tuning parameters $\Delta p_{i}(Q_{0})=p_{i,\mathrm{refH}[Q_{0}]}(Q_{0})-p_{i,\mathrm{cent}}(Q_{0})$ is an estimate for the Apprentice interpolation errors at the cutoff $Q_{0}$ since, for a perfect interpolation and in the absence of statistical errors, we would have $\Delta p_{i}(Q_{0})=0$. This approach also allows to quantify the interpolation uncertainty at the reference cutoff $Q_{0,\mathrm{ref}}$ itself.
To reduce computational cost of this uncertainty estimation method, we carry out this procedure only for the six equidistant cutoff values $Q_{0,m}^{\textrm{H}}\in\{0.75,1.00,...,2.00\}\,$\mathrm{GeV}$$. To obtain an uncertainty estimate for all $Q_{0}$ values we consider, we apply the following averaging procedure. Since the minimization procedure itself is actually not computationally expensive we can evaluate $p_{i,\mathrm{refH}[Q_{0,m}^{\textrm{H}}]}(Q_{0})$ for the $m=1,\ldots,6$ exact Herwig data sets for all $Q_{0}$ values. We then estimate the final interpolation uncertainty of the hadronization parameters by adopting a distance based average,
$\Delta p_{i,\mathrm{inter}}(Q_{0})=\sqrt{\sum_{m=1}^{6}\sum_{n=1}^{26}w(Q_{0},Q^{\prime}_{0,n},Q^{\textrm{H}}_{0,m})\Big{[}p_{i,\mathrm{refH}[Q_{0,m}^{\textrm{H}}]}(Q_{0,n}^{\prime})-p_{i}(Q_{0})\Big{]}^{2}}\,,$ (59)
where the weights are given by
$w(Q_{0},Q^{\prime}_{0,n},Q^{\textrm{H}}_{0,m})\propto e^{-[\sigma_{c}^{-2}(Q^{\textrm{H}}_{0,m}-Q^{\prime}_{0,n})^{2}+\sigma_{s}^{-2}(Q^{\prime}_{0,n}-Q_{0})^{2}]/2},$ (60)
with the sum normalized to one $\sum_{m,n}w(Q_{0},Q^{\prime}_{0,n},Q^{\textrm{H}}_{0,m})=1$. The correlation width $\sigma_{c}=1/32$ GeV in the first exponential ensures that the exact Herwig data tunes at $Q_{0,m}^{\textrm{H}}$ closest to $Q_{0}$ contribute most. The smoothing width $\sigma_{s}=1/16$ GeV in the second exponential ensures that for $Q_{0}$ values in the middle between two $Q^{\textrm{H}}_{0}$ values we obtain about the average of the interpolation uncertainties at these $Q^{\textrm{H}}_{0}$ values.
6.3 Extraction of Migration Matrix Functions
The technical implementation of the migration matrix extraction relies on the true parton level state, in the form of the complete description of all particles in that state, in the HepMC event record for each event that Herwig produces. Before we continue the description we remind the reader about the standard framework that is used to generate observable histograms from MC generator runs: A MC generates each event sequentially. For each event it produces an event record in the HepMC format that also includes all particles in the final state, which are marked by the “final-state” status code (which is just the number 1). Each particle entry $[X,(p_{E},p_{x},p_{y},p_{z})]$ specifies the particle species $X$ and its four-vector $p^{\mu}$. This HepMC event record is read by the analysis framework Rivet. Analyses written for Rivet first carry out a final-state projection which gives the set of all final-state particle four-momenta of the event, $\{p^{\mu}_{i}\}$. This set is then converted to an observable value by an observable projection $\tau=\mathcal{O}(\{p^{\mu}_{i}\})$. For each event this observable value is filled into a histogram to produce the final binned distribution $\mathrm{d}{\sigma}/\mathrm{d}{\tau}$.
In the Herwig code that we use in our analyses we include in addition the true parton level state as defined in Sec. 2.2 in the HepMC event record for each event, with parton level particles marked by a fixed status code, which lies in the value range that a MC can freely use for internal purposes. This allows us to not only do a final-state, i.e. hadron level, projection but also a true parton level one to obtain the parton level four-vectors $\{\hat{p}^{\mu}_{i}\}$, denoted with a hat. This functionality does not only allow for the evaluation of parton level observables, such as $\hat{\tau}=\mathcal{O}(\{\hat{p}^{\mu}_{i}\})$ and $\mathrm{d}{\sigma}/\mathrm{d}\hat{\tau}$, analogous to the hadron level ones, but also for the extraction of the combined set ($(\{\hat{p}^{\mu}_{i}\},\{{p}^{\nu}_{j}\})$ obtained for each event. This provides access to a fully differential probability distribution $P((\{\hat{p}^{\mu}_{i}\},\{{p}^{\nu}_{j}\})$ in the combined parton-hadron level space. We can therefore also extract any parton-hadron level correlation function $\mathrm{corr}[\mathcal{O}_{1}(\{\hat{p}^{\mu}_{i}\}),\mathcal{O}_{2}(\{p^{\nu}_{j}\})]$ or any migration function (or matrix) given by the conditional probability distribution $P(\mathcal{O}(\{p^{\nu}_{j}\})\,|\,\mathcal{O}(\{\hat{p}^{\mu}_{i}\}))$. This allows us to extract the probability of having a hadron level observable value $\tau$ for a given parton level value $\hat{\tau}$. In our phenomenological studies we analyze the behaviour of the migration matrix function of the 2-jettiness observable $(\hat{\tau},\tau)$ and the first moment of the probability distribution $P(\tau\,|\,\hat{\tau})$ in $\tau-\hat{\tau}$. In practice these probability distributions are of course lists or matrices of probabilities as we consider observable bins. There are two approaches to saving the necessary data using Rivet. The first is to fill a very finely binned 2-D histogram in the $(\hat{\tau},\tau)$ variables. The second option is to save the tuple $(w,\hat{\tau},\tau)$ for each event, i.e. unbinned data. The event weight $w$ can in general differ from $1$, in particular for NLO-matched MC simulations. We adopted the second method since it allows us to generate any histograms with arbitrary binning specifications and to calculate any average exactly without the potential need to rerun the MC simulation.
Accounting for the additional dependence on the shower cutoff scale $Q_{0}$ and the hard scattering scale $Q$ in our subsequent analyses, we refer to the probability distribution $P(\tau\,|\,\hat{\tau})$ as $T(\tau,\hat{\tau},\{Q,Q_{0}\}))$, which we already introduced in Sec. 3.3. To be more specific, we concretely study the quantitative behavior of the rescaled transfer functions $\tilde{S}^{\rm MC}(k,\hat{k},\{Q,Q_{0}\})$ and $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ which are derived from $T(\tau,\hat{\tau},\{Q,Q_{0}\}))$ in Eqs. (18) and (19). We remind the reader that $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ is the exact MC analogue of the shape function $S_{\rm had}(\ell,Q_{0})$ appearing in the analytic QCD factorization.
7 Phemomenology of the New Model
Finally, in this section we analyze quantitatively the properties of Herwig‘s default and new dynamic hadronization models from the perspective of (i) the shower cut $Q_{0}$ taking the role of an IR factorization scale and (ii) the $Q_{0}$ scheme dependence of the hadronization migration matrix function $S^{\rm MC}(\ell,\hat{k},\{Q,Q_{0}\})$ in the 2-jettiness dijet region demanded from the QCD factorization, where the hadronization effects are expressed in terms of a shape function. We remind the reader that unless stated otherwise we always use the more precise quartic interpolations for the results that are discussed.
7.1 Shower Cutoff Independence of Hadron Level Observables
Figure 5: Upper panels: Shower cutoff $Q_{0}$-dependence of the Herwig hadron level 2-jettiness distribution in the peak region for the c.m. energies $Q=45$ (left panels), $91.2$ GeV (middle panels) and $Q=200$ GeV (right panels) for the default hadronization model. Lower panels: ratio of the 2-jettiness distributions with respect to the reference $Q_{0,{\rm ref}}=1.25$ GeV tune.
Figure 6: Upper panels: Shower cutoff $Q_{0}$-dependence of the Herwig hadron level 2-jettiness distribution in the peak region for the c.m. energies $Q=45$ (left panels), $91.2$ GeV (middle panels) and $Q=200$ GeV (right panels) for the novel dynamical hadronization model. Lower panels: ratio of the 2-jettiness distributions with respect to the reference $Q_{0,{\rm ref}}=1.25$ GeV tune.
[…]
It is highly instructive to see that this improvement does not only happen for the simulation at $Q=91.2$ GeV, where the tuning procedure is carried out, but also for other hard scattering energies. This is highly rewarding as it shows that the design of the novel dynamical hadronization model properly extrapolates the correct $Q_{0}$ dependence to other energies, that are not controlled directly through the tuning. This property is crucial for the interpretation of the shower cut being an IR factorization scale and the consistency of the hadronization model’s dynamical behavior in the context of QCD.
We have checked that these features are not only realized for 2-jettiness and event-shapes, on which most of our analytic insights concerning the NLL precision of Herwig’s angular ordered parton shower and of the partonic shower cut $Q_{0}$ dependence are based on, but actually for all jet and event-shape related observables that have been measured in $e^{+}e^{-}$ collisions in the past. In Figs. 16 and 17 shown in App. A this is demonstrated displaying the simulations results obtained for the default and the novel dynamical hadronization model, respectively, again for the shower cutoff values $Q_{0}=1$, $1.25$, $1.5$ and $1.75$ GeV for a number of selected observables for in comparison with actual LEP data gathered at $Q=44$, $Q=91$ and $Q=133$ GeV. We have displayed the results for a number of other event-shapes (thrust, heavy jet mass), jet resolution rates as well as charged particle multiplicities measured at the JADE, ALEPH, DELPHI and OPAL experiments MovillaFernandez:1997fr ; JADE:1999zar ; ALEPH:2003obs ; ALEPH:1996oqp ; ALEPH:1991ldi ; DELPHI:1996sen The respective Rivet analyses codes used for the generation of each of the histograms is quoted at the bottom of each panel. The lower sections of the individual panels show ratios with respect to the MC reference simulation results for $Q_{0,\rm{ref}}=1.25$ GeV. The improved shower cutoff $Q_{0}$-independence for the dynamical model in comparison to the default model for all observables is clearly visible. We emphasize again that we have checked that the results shown in Figs. 16 and 17 are representative for all available $e^{+}e^{-}$ Rivet analyses. We also stress that, overall, the quality of the description of the data provided by all the tunes we obtained for the novel dynamical hadronization model is the same as that of the standard Herwig release tune.
7.2 Tuning Quality and Cutoff Dependence of the Tuning Parameters
Figure 7: Minimal GOF values normalized to the sum of squared weights $(\sum_{k}{w_{k}^{2}})$ obtained from the tuning procedures for the default (blue dots) and the novel dynamical hadronization model (orange dots) as a function of the shower cutoff $Q_{0}$. The left panel shows the results for the cubic (o3) and right panel the results for the quartic (o4) interpolations. The colored triangles represent the normalized minimal GOF values obtained from using actual LEP data as the reference date, once for treating $Q_{0}$ as a floating fit parameter and once for using the reference value $Q_{0,\rm{ref}}=1.25$ GeV.
[…]
At this point an interesting aspect to be discussed is to which extent the reference data used for the tuning analyses, which are generated by the MC for the cutoff value $Q_{0,\mathrm{ref}}=1.25$ GeV, represent a good proxy for the real LEP data. Furthermore, it should be also clarified whether the experimental data indeed prefers shower cut values in the range where perturbation theory is still valid, i.e. in the range between $1$ and $2$ GeV that we consider in our discussion. To address these questions we carry out two additional tuning analyses where instead of the MC generated reference data the LEP data is used that also enters the regular Herwig release tunes. In one tuning analysis $Q_{0}$ is treated as a freely floating tuning parameter and in the other the shower cutoff is fixed to $Q_{0,\mathrm{ref}}=1.25$ GeV. The outcome is shown as the colored triangles for the cubic as well as for the quartic interpolations. The minimal GOF function values for these tuning fits are of course not zero since the real data always differs from simulations. For the cubic interpolation both tunes yields practically identical values, with minimal GOF values compatible with the smaller dynamical model GOF values for $Q_{0}>1.4$ GeV. For the quartic interpolation the minimal GOF values for the tune with a freely floating shower cutoff yields $Q_{0}$ values close to unity, and the minimal GOF values are in the range from $4.2$ to $4.6$ for the default and from $3.57$ to $3.64$ for the dynamical model. The results confirm that the MC reference data of our shower cutoff dependent tuning analysis is sufficiently close to the experimental LEP data and, more importantly, that the interval of $1$ to $2$ GeV, where the shower cut $Q_{0}$ can be considered as an infrared factorization scale, is well compatible with the experimental data.
Figure 8: Dependence of the tuned parameters on the shower cutoff $Q_{0}$ for the default hadronization model. The dots represent the central values and the vertical lines the combines statistical and interpolation uncertainty. The orange results are based on cubic and the blue results on the quartic interpolations.
Figure 9: Dependence of the tuned parameters on the shower cutoff $Q_{0}$ for the novel dynamical hadronization model. The dots represent the central values and the vertical lines the combines statistical and interpolation uncertainty. The orange results are based on cubic and the blue results on the quartic interpolations.
Let is now have a look at the $Q_{0}$-dependence of the six hadronization model parameters which are treated as floating parameters in our tuning analyses in each of the hadronization models. As we have already mentioned in Sec. 6.2, for the dynamical model a proper matching to the parton shower (which allows for the shower cutoff $Q_{0}$ to be interpreted as an infrared factorization scale) implies a linear $Q_{0}$-dependence for the “hard” starting scales $\tilde{Q}_{g}$ and $\tilde{Q}_{q}$ of the non-perturbative branching processes and an insensitivity to $Q_{0}$ for the other hadronization parameters. On the other hand, for the default hadronization model (where the shower cutoff is merely another hadronization parameter) similar implications can in principle not necessarily be expected. In Figs. 8 and 9 we show the $Q_{0}$-dependence of the six tuning parameters which were treated as floating parameters for the default and the dynamic models, respectively. We show the results based on the cubic (orange) and the quartic (blue) interpolations. The dots show the central values ${p_{i,\mathrm{cent}}(Q_{0})}$ and the vertical lines stand for the uncertainties which are obtained from the quadratic sum of the statistical und interpolation uncertainty $\Delta p_{i}=\sqrt{(\Delta p_{i,\mathrm{stat}}(Q_{0}))^{2}+(\Delta p_{i,\mathrm{inter}}(Q_{0}))^{2}}$, as described in Sec. 6.2. Overall we see that that tuning results based on the cubic and quartic interpolations are nicely compatible, even though the error bars do not always overlap.
In Fig. 9 we see that, indeed, the parameters $\tilde{Q}_{g}$ and $\tilde{Q}_{q}$ exhibit a $Q_{0}$-dependence that is quite close to linear. At the same time, in comparison the other four hadronization parameters (${\mathtt{Cl}_{\mathtt{max}}}$, ${\mathtt{Cl}_{\mathtt{pow}}}$, $\mathtt{PwtSquark}$ and $\mathtt{PwtDIquark}$) are, indeed, rather $Q_{0}$-insensitive. The behavior is again not perfect, with ${\mathtt{Cl}_{\mathtt{max}}}$ and ${\mathtt{Cl}_{\mathtt{pow}}}$ exhibiting rather large uncertainties, but it can nevertheless be observed rather clearly. In contrast, we can see a somewhat larger $Q_{0}$-dependence of the same four parameters for the default model in Fig. 8. This is visible most prominently for $\mathtt{PwtDIquark}$. Since these parameters directly affect the Baryon production and thus the charge particle multiplicities generated in the simulation, we can conclude that in the default model the shower cutoff value $Q_{0}$ affects these multiplicities so that the tuned value of $\mathtt{PwtDIquark}$ needs to compensate in a significant way. In the novel dynamical hadronization model this rather unnatural feature is not visible. Overall, the results shown in Figs. 8 and 9 again confirm that the novel dynamical hadronization model achieves a better separation of the low-scale hadronization processes from the parton level description provided by the parton shower. This separation is essentially the MC simulation analogue of factorization in analytic QCD studies and an important prerequisite for the parton shower cutoff to be interpreted as a factorization scale.
7.3 Rescaled Migration Matrix Function
Figure 10: Rescaled parton-to-hadron migration matrix funtions for the default hadronization model for the 2-jettiness distribution for the shower cutoff $Q_{0}=1.25$ GeV extracted for c.m. energy $Q=91.2$ GeV. The left panel shows $\tilde{S}^{\rm MC}(k,\hat{k},\{Q,Q_{0}\})$ and the right panel $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$, which is analogous to the shape function.
Figure 11: Rescaled parton-to-hadron migration matrix funtions for the novel dynamical hadronization model for the 2-jettiness distribution for the shower cutoff $Q_{0}=1.25$ GeV extracted for c.m. energy $Q=91.2$ GeV. The left panel shows $\tilde{S}^{\rm MC}(k,\hat{k},\{Q,Q_{0}\})$ and the right panel $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$, which is analogous to the shape function.
Let us now start the discussion on the MC results for the rescaled migration matrix functions $\tilde{S}^{\rm MC}(k,\hat{k},\{Q,Q_{0}\})$ and $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ defined in Eqs. (18) and (19), respectively, for the default and the dynamic hadronization models obtained from the 2-jettiness $\tau$ distributions. For a given (true) parton level soft momentum $\hat{k}=Q\hat{\tau}$ and shower cutoff $Q_{0}$, the migration function $\tilde{S}^{\rm MC}(k,\hat{k},\{Q,Q_{0}\})$ gives the distribution of hadron level momenta $k=Q\tau$, while $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ gives the distribution of the non-perturbative soft momenta $\ell=k-\hat{k}$ that the hadronization adds to the parton level configuration. We remind the reader that $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ is the MC analogue of the infrared cutoff-dependent shape function $S_{\rm had}(\ell,Q_{0})$ in the dijet 2-jettiness QCD factorization formula discussed in Sec. 3.2. Details concerning the extraction of the migration matrix functions from the Herwig simulations have been explained in Sec. 6.3.
In the left panel of Fig. 10 the MC migration function $\tilde{S}^{\rm MC}(k,\hat{k},\{Q,Q_{0}\})$ is shown for the Herwig default cluster hadronization model. It has been extracted at the c.m. energy $Q=91.2$ GeV for the reference $Q_{0}=1.25$ GeV tune (explained in more detail in Sec. 6.1) in the range $0\leq k,\hat{k}\leq 8$ GeV. For $Q=91.2$ GeV this corresponds to thrust values between 00 and $0.088$. The corresponding shifted migration function $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ is shown in the right panel. In Fig. 11 the analogous migration functions are shown for the novel dynamical model. The shape and structures shown in both figures are representative for the migration matrix functions for all cases we obtain in our analyses
We see that for $\hat{k}>2$ GeV the shape of the migration matrix functions for both hadronization models are very well consistent with the expectations from a shape function, which states that the migration function $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ should be independent of the partonic momentum $\hat{k}$. We can spot deviations from this expectation for $\hat{k}<2$ GeV for the default as well as for the novel dynamical hadronization model. While for the default hadronization model the migration function $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ for $\hat{k}<2$ GeV is considerably flatter than for $\hat{k}>2$ GeV, for the dynamic hadronization model it is much peakier. These features play an essential role in the more important quantitative analyses we carry out in Sec. 7.4.
Before moving on, let us comment on a general feature of the migration functions. It concerns that $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ is finite for $\ell$ values down to $-1$ GeV. We can see this feature clearly in the right panels of Figs. 10 and Fig. 11 for partonic values $\hat{k}>1$ GeV. Even though the major portion of $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ is located at positiv $\ell$, which means that the bulk of hadronization corrections shift the thrust distribution towards larger thrust values, this negative tail shows that hadronization can sometimes also decrease the thrust value. Since this feature arises from the tuning to actual LEP data, it means that at least for the Herwig simulation implementation this particular feature of the hadronization corrections is demanded by data. For $\hat{k}\geq 1.5$ GeV this behaviour arises consistently and becomes independent of the value of $\hat{k}$ also for other $Q$ and $Q_{0}$. This feature is per se not at all problematic and also consistent with a shape function within QCD factorization. However, for partonic $\hat{k}$ values below $1.5$ GeV the migration function is not capable to build up such a negative tail, simply because the physical thrust values are restricted to be positive. This entails that the first $\ell$-moment of the rescaled migration function $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ (see Eq. (13)) for $\hat{k}$ values below $1.5$ GeV is always larger than for values larger than $2$ GeV, where the first moment stabilizes. This feature is absent in the shape function $S_{\rm had}(\ell,Q_{0})$, which is strictly $\hat{k}$-independent. We come back to this feature in our discussion below.
7.4 Shower Cutoff and Energy Dependence of the Migration Matrix Function
Figure 12: Rescaled parton-to-hadron migration matrix function $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ as a function of $\ell$ for the default hadronization model for $Q_{0}=1.25$ and a number of different partonic $\hat{k}=\hat{\tau}/Q$ soft momenta. The upper panels show the results for some $\hat{k}<2$ GeV and the lower panels for larger $\hat{k}$ values. The left, middle and right panels have been extracted from the 2-jettiness distributions for c.m. energies $Q=45$, $91.2$ and $200$ GeV, respectively.
Figure 13: Rescaled parton-to-hadron migration matrix function $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ as a function of $\ell$ for the novel dynamical hadronization model for $Q_{0}=1.25$ and a number of different partonic $\hat{k}=\hat{\tau}/Q$ soft momenta. The upper panels show the results for some $\hat{k}<2$ GeV and the lower panels for larger $\hat{k}$ values. The left, middle and right panels have been extracted from the 2-jettiness distributions for c.m. energies $Q=45$, $91.2$ and $200$ GeV, respectively.
Finally, let us now discuss at the more quantitative level the properties of the migration matrix functions $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ we have obtained from our $Q_{0}$-dependent tuning analyses for the Herwig default and the novel dynamical hadronization model.
[…]
Figure 14: First moment $\Omega_{1}^{\rm MC}(\hat{k},Q,Q_{0})$ of the rescaled parton-to-hadron migration function over $\hat{k}$ for the default (left panel) and the novel dynamical hadronization model (right panel) extracted from the 2-jettiness distribution simulated for shower cutoff $Q_{0}=1.25$ GeV at c.m. energies $Q=45$ GeV (blue), $91.2$ (orange) and $200$ GeV (green dots).
To gain a more quantitative insight into the $Q_{0}$-, $\hat{k}$- and $Q$-dependence of the migration function $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ it is instructive to have a close look on its first moment defined by
$\Omega_{1}^{\rm MC}(\hat{k},Q,Q_{0})\,\equiv\,\frac{1}{2}\int\!\mathrm{d}\ell\,\ell\,S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})\,,$ (61)
in analogy to the QCD factorization’s shape function first moment $\Omega_{1}(Q_{0})$ given in Eq. (13). The latter only depends on the shower cut $Q_{0}$, but is independent of $\hat{k}$ or $Q$. In Fig. 14 the value of $\Omega_{1}^{\rm MC}(\hat{k},Q,Q_{0})$ is displayed for $30<\hat{k}<$ 3 GeV and $Q=45$ GeV (blue), $91.2$ GeV (orange) and $200$ GeV (green) for $Q_{0}=1.25$ GeV. The left panel shows $\Omega_{1}^{\rm MC}$ for the default hadronization model and the right panel that for the novel dynamical model. For both we only see a mild $Q$- and $\hat{k}$-dependence for $\hat{k}\gtrsim 1.5$ GeV, where the dependence is slightly smaller for the dynamical model. For small $\hat{k}<1.5$ GeV, however, the dependence on $\hat{k}$ and $Q$ is enormous for the default model. For the smallest $\hat{k}$ value we have $\Omega_{1}^{\rm MC}\approx 2.1$ GeV for all $Q$ values for the dynamical model. On the other hand, $\Omega_{1}^{\rm MC}$ varies wildly with $Q$ and even reaches $5.5$ GeV for $Q=200$ GeV. It is this uncontrolled behavior for the default hadronization model that is responsible for the stronger $Q_{0}$ dependence visible in Fig. 5 in comparison to Fig. 6.
Nevertheless, also for the dynamic model $\Omega_{1}^{\rm MC}(\hat{k},Q,Q_{0})$ still exhibits a visible $\hat{k}$ dependence for $\hat{k}<1.5$ GeV which is in principle not compatible with the shape functions first moment. This is again related to the fact already mentioned at the end of Sec. 3.2, namely that for the Herwig MC setup the hadronization effects can sometimes lead to a decrease of the hadron level 2-jettiness value $\tau$ with respect to its parton level value $\hat{\tau}$. For values $\hat{k}\lesssim 1$ GeV this behavior cannot be realized any longer because the physical 2-jettiness values cannot become negative. Since the hadronization model’s capability to increase the hadron level 2-jettiness value remains unchanged for these low $\hat{k}$ values, this feature leads to the more peaky shape of the $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ just mentioned in the discussion of Fig. 6 above, which unavoidably leads to an increasing first moment for $\hat{k}\lesssim 1$ GeV.
[…]
$\Omega_{1}(Q_{0})-\Omega_{1}(Q_{0,\mathrm{ref}})\,=\,\frac{1}{2}\,\Delta_{\rm soft}(Q_{0},Q_{0,\mathrm{ref}})\,,$ (62)
where $\Delta_{\rm soft}(Q_{0},Q_{0,\mathrm{ref}})$ is determined from the $R$-evolution equation in Eq. (10).
We see that for $\hat{k}>1$ GeV the moment difference obtained for the dynamical model (lower panels) are indeed nicely compatible with Eq. (62). The visible small discrepancies are related to quadratic and higher order effects in $Q_{0}$ which the linear approximation of the evolution equation does not capture. They are also compatible with the corresponding discrepancies concerning the $Q_{0}$ dependence of the partonic cumulant difference already discussed in Sec. 3.3 for Fig. 1. In stark contrast, for the default model results (upper panels) there is no sign of a similar agreement with Eq. (62) or of any stability with respect to the value of $\hat{\tau}$. We see that for the default model the moment differences are completely unrelated to the shower cut dependence expected from QCD factorization, and thus essentially uncontrolled from the QCD perspective. We stress, that this uncontrolled $Q_{0}$-dependence of the hadronization corrections from the default hadronization model takes place even for large $\hat{k}$ values where the migration function $S^{\rm MC}(\ell,\{\hat{k},Q,Q_{0}\})$ appears stable (also with respect to changes of $Q$) as shown in the lower panels of Fig. 12. There is furthermore no stability concerning the $\hat{k}$ dependence of $\Omega_{1}^{\rm MC}(\hat{k},Q,Q_{0})-\Omega_{1}^{\rm MC}(\hat{k},Q,Q_{0,\mathrm{ref}})$ with respect the hard scattering scale $Q$.
It is instructive to discuss this failure of the default model from a the perspective of overall size of the hadronization corrections. As we can see from Fig. 14, for $Q_{0,\mathrm{ref}}=1.25$ we have $\Omega_{1}^{\rm MC}$ in the range between $1.2$ and $1.5$ GeV for partonic momenta $\hat{k}>2$ GeV. This should be compared to the overall variation of $\Omega_{1}^{\rm MC}$ in the range between $Q_{0}=1.0$ GeV and $Q_{0}=1.75$ GeV predicted by $R$-evolution equation (10) obtained from QCD factorization. The latter amounts to around $0.5$ GeV, as we see in the lower panels of Fig. 15. Howerver, because the choice of reference shower cut $Q_{0,\mathrm{ref}}$ is arbitrary, and we do not have any principle means to tell for which value of $Q_{0}$ the default model may have the best agreement with the expectations from QCD factorization (which can only be tested by a $Q_{0}$-dependence consistent with QCD factorization), this implies that the size of the hadronization corrections to the (true) partonic thrust distribution that is provided by the default model is inconsistent with QCD factorization in the range of $40$%. This is substantially worse in comparison to the new default model, where we have consistency at the level of better than $10$%. In the context of the shower cut $Q_{0}$ adopting the role of an IR factorization scale, so that the hadronization effects have a well-defined and controlled scheme dependence so that they can be assigned field theoretic meaning, the default model thus fails quite badly. In this respect the novel dynamical hadronization model performs substantially better.
Overall, the new dynamic hadronization model provides a significant improvement concerning the control of the shower cut $Q_{0}$ as an IR factorization scale. This feature is a prerequisite to combine the hadronization corrections implemented and quantified in the hadronization model with parton level theoretical calculations in a meaningful and systematic manner.
8 Conclusions and Outlook
In this article we further promote the idea of the parton shower cutoff $Q_{0}$ for MC simulations being an infrared factorization scale that separates, in a controlled manner and compatible with QCD, perturbative and non-perturbative hadronization effects. In this context, features of the MC’s hadronization model may be given a systematic field theoretic meaning and, at the same time, QCD parameters appearing in the parton shower may be related to their renormalization scheme dependent counter parts appearing in analytic QCD computations. An important prerequisite to achieve this is to have a parton shower algorithms that has at least NLL precision for the observable considered. However, the second essential and equally important prerequisite is to have a hadronization model that can properly match the unavoidable infrared cutoff $Q_{0}$-dependence of the parton level description that emerges from the parton shower. This is a highly nontrivial condition on the hadronization model since it entails that the parton shower cutoff $Q_{0}$ is not treated as a tuned hadronization parameter. Rather, the hadron level MC observable description should be equivalent for different $Q_{0}$ values, at least within some low energy interval, where QCD perturbation theory can still be trusted.
In this work we have investigated these features using the angular ordered parton shower and the cluster hadronization model implemented in the Herwig 7.2 MC event generator focusing primarily on the 2-jettiness distribution in $e^{+}e^{-}$ annihilation for which the angular ordered parton shower is NLL precise. In the earlier work Hoang:2018zrp some of us have analyzed the gluon transverse momentum cutoff $Q_{0}$-dependence that emerges from the angular order parton shower for the 2-jettiness distribution in detail at the subleading ${\cal O}(\alpha_{s})$ level using QCD factorization in the dijet region. It was also shown that the Herwig parton shower implementation satisfies the $Q_{0}$ evolution equation from QCD factorization very well. In this work we have now extended our investigations concerning the parton shower infrared cutoff being a factorization scale from the perspective of the Herwig cluster hadronization model.
The default cluster hadronization model has not been designed in a way so that it systematically matches to the parton shower for a range of $Q_{0}$ values. This is related to a number of features in the cluster formation and fission dynamics that yield good data description through the tuning, but are otherwise ad-hoc and not compatible with the processes that happen in the parton shower evolution close to the cutoff $Q_{0}$. As a result, the hadronization effects to the 2-jettiness distribution provided by the default Herwig cluster model do not satisfy the QCD constraints that emerge from $Q_{0}$ being promoted to a factorization scale in a satisfactory manner. We have demonstrated this in a number of tests based on tuning analyses for different $Q_{0}$ shower cutoff values where $Q_{0}$ is treated as an external scale and not a tuned parameter. These tests involve analyses of the $Q_{0}$-dependence of hadron level simulations for 2-jettiness and also other event-shapes and observables and of the parton-to-hadron migration matrix for 2-jettiness for which QCD factorization provides nontrivial constraints.
To improve the cluster hadronization model we have added a number of features to the cluster formation and cluster fission processes that mimic the gluon emission and splitting dynamics that takes place in the parton shower. These modifications provide a clearer separation of model parameters that are expected to be correlated to $Q_{0}$, governing the ’hard’ aspects of the cluster hadronization dynamics, from those which govern ’soft’ hadron formation aspects that should be rather $Q_{0}$-independent. In our $Q_{0}$-dependent tuning studies we found that this novel dynamical cluster hadronization model performs substantially better concerning the $Q_{0}$-invariance of Herwig’s hadron level predictions as well as concerning the QCD factorization constraints on the 2-jettiness parton-to-hadron migration matrix. We emphasize that our analyses also involved hard scattering energies that are not accounted for in the reference data used for the tuning. This shows that the novel dynamical hadronization model properly scales the consistency to QCD factorization to other hard scattering energies.
Even though the novel dynamical hadronization model we have designed in this article is not perfect, its features provide an important step forward in promoting the hadronization corrections encoded in MC generators to have a well-defined scheme in the QCD context, as also discussed previously in Ref. Platzer:2022jny . This is an essential aspect that should be followed in parallel to the ongoing developments of subleading order precise parton shower algorithms such that the parameters of the parton shower as well as the hadronization model can acquire a systematic QCD field theoretic meaning. Beyond a more precise and consistent description of experimental data, important potential applications of such improvements are the determination of QCD parameters directly from MC studies as well as a systematic quantification of hadronization corrections to analytic QCD calculations from MC simulations. In an upcoming article we will investigate the former application from the perspective of the MC top quark mass parameter.
Acknowledgments
We acknowledge support by the FWF Austrian Science Fund under the Project No. P32383-N27 and under the FWF Austrian Science Fund Doctoral Program “Particles and Interactions” No. W1252-N27 as well as partial support by the COST Action No. CA16201 Particleface. This work has also been supported in part by the European Union’s Horizon 2020 research and innovation programme as part of the Marie Skłodowska-Curie Innovative Training Network MCnetITN3 (grant agreement No. 722104). We are grateful the Erwin-Schrödinger International Institute for Mathematics and Physics for partial support during the Thematic Programme “Quantum Field Theory at the Frontiers of the Strong Interactions”, July 31 - September 1, 2023.
Appendix A Comparison of Simulations with LEP Data
Figure 16: Selected rivet analyses for $e^{+}e^{-}$ data (black) versus Herwig simulations for the default hadronization model for shower cutoff values $Q_{0}=1$ (blue), $1.25$ (orange), $1.5$ (green) and $1.75$ GeV (red). The ratios in the lower panel sections are shown w.r. to the simulations for $Q_{0}=1.25$ GeV.
Figure 17: Selected rivet analyses for $e^{+}e^{-}$ data (black) versus Herwig simulations for the novel dynamical hadronization model for shower cutoff values $Q_{0}=1$ (blue), $1.25$ (orange), $1.5$ (green) and $1.75$ GeV (red). The ratios in the lower panel sections are shown w.r. to the simulations for $Q_{0}=1.25$ GeV.